### v2

# Optimizing Information Access in Networks via Edge Augmentation

Abstract
Given a graph $G=(V,E)$ and a model of information flow on that network, a fundamental question is to understand whether all nodes have sufficient access to information generated at other nodes in the graph. If not, we can ask if a small set of “interventions” in the form of edge additions improve information access. Formally, the broadcast value of a network is defined to be the minimum over pairs $u,v\in V$ of the probability that an information cascade starting at $u$ reaches $v$. Having a high broadcast value ensures that every node has sufficient access to information spreading in a network, thus quantifying fairness of access. In this paper, we formally study the Broadcast Improvement problem: given $G$ and a parameter $k$, the goal is to find the best set of $k$ edges to add to $G$ in order to maximize the broadcast value of the resulting graph.
We develop efficient approximation algorithms for this problem. If the optimal solution adds $k$ edges and achieves a broadcast of $\beta^{*}$, we develop algorithms that can (a) add $k$ edges and achieve a broadcast value roughly $(\beta^{*})^{4}/16^{k}$, or (b) add $O(k\log n)$ edges and achieve a broadcast roughly $\beta^{*}$. We also provide other trade-offs that can be better depending on the parameter values. Our algorithms rely on novel probabilistic tools to reason about the existence of paths in edge-sampled graphs, and extend to a single-source variant of the problem, where we obtain analogous algorithmic results. We complement our results by proving that unless P=NP, any algorithm that adds $O(k)$ edges must lose significantly in the approximation of $\beta^{*}$, resolving an open question from prior work.
1 Introduction
Social networks are one of the primary sources of information in modern society. From news and current events, to job opportunities, to public health announcements, information flows between users in social networks, often more effectively than other channels. However, such modes of flow come with inherent drawbacks: well-connected and “influential” groups of nodes in the network (or the communication graph) can have an unfair advantage in how quickly they gain access to the information, or in controlling the flow of information. This has led to a significant body of research around fairness of information access on social networks [16, 31, 28, 1, 22, 3, 4, 31, 29, 23, 26].
Within computer science, the classic work of Kempe, Kleinberg, and Tardos [25] formalized the independent cascade model, which has become the standard way of modeling information spread in networks. While originally defined via an iterative spreading process, the probability of two nodes sharing information under Independent Cascade (which we refer to as the access probability or the proximity of the two nodes; see Section 2 for formal definitions) is equivalent to the probability that they are in the same connected component in a graph generated by adding each edge in the initial graph independently with probability $\alpha$.
Starting with [25], several works have studied how to optimize information spread in networks. A large body of work focuses on influence maximization, which is the problem of finding an optimal “seed set” of nodes (those initially possessing the information) that maximize the expected number of nodes that subsequently receive the information, under the independent cascade model. While this is ideal for applications like advertising where the goal is to maximize the spread, it may result in an asymmetry of access as certain sub-groups of nodes may be isolated. Thus, a recent line of work [2, 22, 30] initiated the study of broadcast, defined as the minimum proximity between any two nodes in the graph. A high broadcast value ensures that information originating at any node has a high probability of reaching every other node in the graph. In applications like job postings and public announcements, the works above argue that broadcast is a better objective from a fairness perspective.
Improving broadcast. Note that broadcast is a structural property of the graph itself, and treats information from (and to) all the nodes as equally important. In this sense, broadcast is similar in spirit to the diameter of a graph (a connection that our results leverage), however, due to its definition in terms of information propagation probabilities, its value also depends on the number of paths, not just the length. [^1] Consider, for instance, a graph $(V,E)$ in which $V$ is partitioned into $V_{1},V_{2},\dots,V_{n/\log^{2}n}$, where $|V_{i}|=\log^{2}n$ for all $i$, and the edges consist of complete graphs (a) within $V_{i}$, and (b) between $V_{i}$ and $V_{i+1}$, for all $i\geq 1$. In this case, the diameter is $n/\log^{2}n$, whereas the broadcast value (assuming the propagation parameter $\alpha=0.5$) will be $\Theta(1)$. We study the problem of augmenting a graph to increase the broadcast value. Given a budget $k$, we study Broadcast Improvement, the problem of adding $k$ edges to the graph so as to maximize the broadcast value of the resulting graph. The work of [2] designed and emipirically evaluated several heuristics for Broadcast Improvement, without providing any theoretical guarantees. [^2] Bashardoust et al. [2] call this problem MaxWelfare-Augmentation. They also leave open the question of determining if Broadcast Improvement is NP-hard. In this paper, we give a theoretical framework for broadcast improvement and address both these questions.
At a high level, Broadcast Improvement is related to network design, an area that is studied in both the data mining and the theoretical computer science communities. Especially relevant to our work are results on edge augmentations that improve various connectivity metrics associated with nodes (see, e.g., [6, 27, 9]). From a techniques perspective, we also draw insights from prior algorithms for Diameter Reduction [11, 7, 17], which can be viewed as the analogue of our problem when the information access between two vertices is modeled as shortest path distance. This is also part of the larger literature on parameterized complexity of edge modification to achieve various graph properties; we refer the reader to the recent survey by Crespelle et. al. for an overview [10].
1.1 Our Results
We provide algorithms and lower bounds for the Broadcast Improvement problem, where we are given a parameter $k$ and the goal is to compete with the optimal broadcast possible by adding $k$ edges to $G$. Denoting this optimal broadcast value by $\beta^{*}$, we give bicriteria algorithms that trade off between the number of edges added and the obtained broadcast value. In our algorithms, we also take into account the dependence on the parameter $\alpha$, the edge activation parameter (probability with which information spreads across a single edge). This is because depending on the application and network structure, information graphs have been used with a range of values for $\alpha$, from around $0.01$ [5] to $0.6$ [2]. We obtain algorithms with different trade-offs between $\alpha$ and $1/k$; see Table 1 for a summary.
Table 1: Summary of algorithms, assuming optimal broadcast on adding $k$ edges is $\beta^{*}$.
Algorithm Number of Edges Added Broadcast Guarantee
$k$-center I $12k-$ 1 $\frac{(\beta^{*})^{4}\alpha^{2}}{(1+2k\alpha)^{4}}$
$k$-center II $k$ $\frac{(\beta^{*})^{4}\alpha^{2}}{16^{k+1}}$
Constant witness I $O(k^{3}\log n)$ $\Omega\left(\frac{\beta^{*}}{k^{4}}\right)$
Constant witness II $O(k^{2}\log n)$ $\Omega\left(\frac{\beta^{*}\alpha}{k^{2}}\right)$
Submodularity-based $O(\frac{k}{\varepsilon}\log n)$ $\Omega\left(\frac{\beta^{*}\alpha^{2+\varepsilon}}{k^{2}}\right)$ ($\forall\varepsilon>0$)
Algorithmic Results. We obtain two kinds of approximation guarantees for Broadcast Improvement, polynomial approximations, where we guarantee that the broadcast value of the resulting graph is $(\beta^{*})^{O(1)}f(k,\alpha)$ for appropriate $f()$, and linear approximations, where we guarantee a broadcast value of the form $\beta^{*}\cdot f(k,\alpha)$. Our contributions are as follows.
  • By observing that the negative logarithms of proximities form a metric, we show a connection between Broadcast Improvement and the classic Metric $k$-Center problem. This yields a simple algorithm for Broadcast Improvement that adds $2k$ edges and achieves broadcast roughly $(\beta^{*})^{4}/k^{4}$. We then refine this argument to obtain an algorithm that adds precisely $k$ edges, while achieving broadcast roughly $(\beta^{*})^{4}/16^{k}$.
  • In order to obtain linear approximations on $\beta^{*}$, we give two approaches. First, we show that there exist near-optimal solutions with a special structure (a framework that we call constant witnesses). Using this, we give algorithms that obtain linear approximations whilst adding $O(k^{2}\log n)$ and $O(k^{3}\log n)$ edges respectively. (See Table 1 and Section 4.) Second, we give an algorithm based on submodularity. Even though the pairwise access probability as a function of added edges is not submodular (1), we argue that for subsets of edges forming a star, access to the center of the star possesses a submodularity property! We leverage this to obtain a solution that adds $O(k\log n)$ edges, and guarantees broadcast roughly $\frac{\beta^{*}\alpha^{2}}{k^{2}}$ (see Section 5).
Our algorithms can also be extended (a) to obtain improved guarantees under an appropriate doubling dimension assumption on the metric defined by the logarithm of the proximities (see Section 3.3), and (b) to the single source version of the problem, which we call Reach Improvement (see Section 7).
Hardness of Broadcast Improvement. Complementing our algorithmic results, we prove an inapproximability result for Broadcast Improvement. Note that even NP-hardness was left as an open problem in [2]. We prove that unless P=NP, no efficient algorithm that outputs $O(k)$ edges can guarantee a broadcast better than $(\beta^{*})^{6/5-\varepsilon}$. This hardness result also suggests that the $\log n$ overhead in our linear approximations to broadcast might be necessary.
2 Preliminaries
We use standard graph-theoretic definitions and notations throughout. Given a graph $G=(V,E)$ and a set $S\subseteq V^{2}\setminus E$, we write $G+S=(V,E\cup S)$ to denote the graph formed by adding the edges $S$ to $G$. For a positive integer $n$, we use the notation $[n]=\{1,2,\ldots,n\}$ for the set of integers from $1$ to $n$. Unless otherwise specified, all logarithms have base $2$. When discussing a probability event $\mathcal{E}$, we use the notation $\bar{\mathcal{E}}$ to indicate the complement of $\mathcal{E}$.
An information graph $G=(V,E,\alpha)$ is a connected graph together with a sampling probability $\alpha$. Throughout, we will be interested in graphs formed by deleting each edge in an information graph $G$ independently with probability $(1-\alpha)$. Alternatively, one may think of each edge as being “activated” (spreading information) with probability $\alpha$ as in Independent Cascade [25]. We call an object formed via these processes a graph sampled from $G$, or simply a sampled graph when the context is clear. For any pair of vertices $u,v$ in an information graph $G$, the probability with which $u$ and $v$ are connected in a sampled graph is the proximity (or access) of $u$ to $v$ in $G$, denoted $\mathcal{P}_{G}({u},{v})$.
We will often discuss proximity in terms of paths. A path $p$ from $u$ to $w$ of length $\ell$ is an alternating sequence of distinct vertices and edges, $u=v_{0},e_{1},v_{1},e_{2},\ldots e_{\ell},v_{\ell}=w$, so that $v_{i}$ is a shared endpoint of $e_{i}$ and $e_{i+1}$ for $i\in[\ell-1]$. In the context of the path $p$, we say that $v_{i-1}$ is the leading vertex of the edge $e_{i}$, and that $v_{i}$ is the trailing vertex of $e_{i}$. If $j>i$, we say that $v_{i}$ precedes $v_{j}$ on $p$. The segment $p[v_{i},v_{j}]$ is the portion of $p$ beginning with $v_{i}$ and ending with $v_{j}$. Given a set of paths $P$ on which $u$ always precedes $v$, we extend the definitions of precedes, segment, and the notation $P[u,v]$ in the natural way, i.e., we say that $u$ precedes $v$ on $P$ and that the segment $P[u,v]$ is the set of segments $p[u,v]$ for all $p\in P$. The contribution of a path $p$ is the probability with which $p$ exists in a sampled graph. Similarly, the contribution of a set of paths $P$ is the probability with which at least one path in $P$ exists in a sampled graph. At times, we use the notation $\Pr[P]$ for the contribution of a set of paths $P$. We observe that for any two vertices $u$ and $v$, $\mathcal{P}_{G}({u},{v})$ is equal to the contribution of the set $P_{uv}$ of all paths from $u$ to $v$ in $G$.
We note that proximity is a symmetric function, i.e., for all pairs of vertices $u$ and $v$, $\mathcal{P}_{G}({u},{v})=\mathcal{P}_{G}({v},{u})$. Furthermore, for every vertex $v$, $\mathcal{P}_{G}({v},{v})=1$. We define the reach of vertex $v$, denoted $\Upsilon_{G}({v})$, as the minimum across all vertices $u\neq v$ of $\mathcal{P}_{G}({v},{u})$. The broadcast of an information graph $G$, denoted $\beta({G})$, is the minimum across all vertices $v$ of $\Upsilon_{G}({v})$. Equivalently, $\beta({G})$ is the minimum pairwise proximity across all pairs of vertices in $G$. We can now define the problem we study:
Broadcast Improvement
Input: An information graph $G=(V,E,\alpha)$ and a non-negative integer $k$. Task: Add a set $F$ of $k$ edges to $G$ such that the broadcast of $(V,E\cup F,\alpha)$ is maximized.
Given an instance $(G,k)$ of Broadcast Improvement, we write $\beta^{*}$ for the optimum broadcast achievable by adding $k$ edges to $G$. We note that computing exact proximity values is #P-hard [8], but standard Monte Carlo approaches have been employed to estimate them, e.g., as seen in [2]. We proceed with the assumption that we possess an oracle which can compute the proximity of any two vertices in polynomial time.
3 Polynomial Approximations Based on Metric $k$-Center
In Section 3.1 we illustrate a connection between Broadcast Improvement and Metric $k$-Center, and use it to obtain a simple bicriteria approximation. Next, in Section 3.2 we show that with more work we can avoid any violation of the edge budget. Finally, in Section 3.3 we mention that our broadcast guarantees can be improved in running time which is fixed-parameter tractable in the combination of $k$ and the doubling dimension of a particular metric space associated with information graphs.
3.1 A Bicriteria Approach
We begin this section with a pair of structural observations. The first is that the negative logarithm of proximity is a metric on the vertices of an information graph. Formally, let $G=(V,E,\alpha)$, and define $\phi\colon V^{2}\rightarrow\mathbb{R}$ by $\phi(u,v)=-\log\mathcal{P}_{G}({u},{v})$. We claim that $(V,\phi)$ is a metric space. Symmetry and non-negativity are trivial, and every point has distance zero to itself because $\mathcal{P}({v},{v})=1$ for all $v\in V$. The triangle inequality follows quickly from the observation that for all $u,v,w\in V,\mathcal{P}({u},{v})\geq\mathcal{P}({u},{w})\cdot\mathcal{P}({w},{v})$. Henceforth, we refer to $(V,\phi)$ as the implied metric of $G$.
The second is a technical lemma which formalizes the intuitive notion that adding $k$ edges “far” from a vertex $v$ cannot improve the reach of $v$ by very much. Formally, given a vertex $v$ and $x\geq 0$, the $x$-neighborhood of $v$, denoted $N_{x}[{v}]$, is the set $\{u\in V\colon\mathcal{P}_{G}({v},{u})\geq x\}$. In other words, $N_{x}[{v}]$ is the set of vertices which have probability at least $x$ of getting information from $v$. The main idea of the proof is to impose an appropriate equivalence relation on the set of all paths between $v$ and $u$, and then consider the equivalence classes separately. This technique will appear repeatedly throughout the work, often in more complicated forms.
Lemma 1.
Consider $x\geq 0$ and two vertices $v,u$ in an information graph $G$, such that $\mathcal{P}_{G}({v},{u})<x$. Suppose that an augmented information graph $G^{\prime}$ is formed by adding $k$ edges whose endpoints are disjoint from $N_{x}[{v}]$. Then $\mathcal{P}_{G^{\prime}}({v},{u})\leq x(1+2k\alpha)$.
Proof.
Let $u\notin N_{x}[{v}]$ be any vertex with $\mathcal{P}_{G}({u},{v})<x$. We consider all paths from $v$ to $u$ in $G^{\prime}$ which include at least one newly added edge. We partition these paths according to the first new edge to appear along the path. For each of the $k$ equivalence classes, we partition again according to which endpoint is the leading vertex of the first new edge. This procedure produces at most $2k$ equivalence classes in total. We now bound the contribution of each equivalence class. Consider the equivalence class defined by new edge $wz$, with $w$ being the leading vertex. Every path in this class begins with a segment from $v$ to $w$ which exists in $G$, and then extends via the edge $wz$. Consequently, the contribution of this set of paths is at most $\mathcal{P}_{G}({v},{w})\alpha\leq x\alpha$. The result follows from applying the union bound on each of our $2k$ equivalence classes, plus the contribution of paths which use no new edges.
We note that the bound given by Lemma 1 is asymptotically tight; see the example given in Figure 1. In this case, for sufficiently small $\alpha$ and $x$, we have
$\mathcal{P}_{G^{\prime}}({v},{u})\approx 1-(1-x\alpha)^{k}(1-x)\approx 1-(1-kx\alpha)(1-x)=x+\Omega(xk\alpha).$
$v$$q_{1}$$q_{2}$$q_{k}$u
Figure 1: An example demonstrating the tightness of the bound given in Lemma 1. The dotted circle represents the $x$-neighborhood of $v$. Each of $u,q_{1},q_{2},\ldots q_{k}$ has proximity $x-\varepsilon$ to $v$, for some small $\varepsilon$. In the augmented information graph $G^{\prime}$, $k$ (dashed) edges are added between $u$ and each of $q_{1},q_{2},\ldots q_{k}$.
We are now ready to present the main ideas of this section. We begin by deriving an upper bound for Broadcast Improvement in terms of an approximation for Metric $k$-Center, for which the input is a metric space $(X,d)$ and a non-negative integer $k$, and the task is to choose $k$ points such that $\max_{v\in X}\min_{s\in S}d(v,s)$ (the resulting radius) is minimized.
Lemma 2.
Let $(G=(V,E,\alpha),k)$ be an instance of Broadcast Improvement, $(V,\phi)$ be the implied metric of $G$, $c\geq 1$, and $x$ be a value such that $-\log x$ is at most a $c$-factor larger than the optimum radius for Metric $2k$-Center on the instance $((V,\phi),2k)$. Then $\beta^{*}\leq x^{1/c}(1+2k\alpha)$.
Proof.
We first claim that for every vertex set $S\subseteq V$ of size at most $2k$, there exists some vertex $u\in V$ with the property that $\max_{s\in S}\mathcal{P}_{G}({u},{s})\leq x^{1/c}$. Otherwise, every vertex has proximity greater than $x^{1/c}$ to some member of $S$, meaning that every vertex has distance less than $-\frac{1}{c}\log x$ to some member of $S$ in the metric space $(V,\phi)$, contradicting our definition of $x$. Now, let $A\subseteq V$ be the endpoints of the edges added in some optimum solution to Broadcast Improvement. Then $A$ has size at most $2k$. Let $u$ be a vertex such that $\max_{v\in A}\mathcal{P}_{G}({u},{v})\leq x^{1/c}$. Equivalently, $N_{x^{1/c}}[{u}]$ is disjoint from $A$. Then by Lemma 1, $\beta^{*}\leq\Upsilon_{G^{*}}({u})\leq x^{1/c}(1+2k\alpha)$. ∎
We convert Lemma 2 into an algorithm for Broadcast Improvement by using a standard 2-approximation, e.g., as given by Gonzalez [18] or Hochbaum and Shmoys [20] to select $2k$ centers, and then adding edges such that these centers form a star.
Theorem 1.
There exists a polynomial-time algorithm which produces an information graph with broadcast at least $\frac{(\beta^{*})^{4}\alpha^{2}}{(1+2k\alpha)^{4}}$ by adding at most $12k-$ 1 edges.
Proof.
We begin by approximating Metric $k$-Center on the implied metric $(V,\phi)$ of $G$ with parameter $2k$. In particular, we obtain a set $S\subseteq V$ of size $2k$ and a value $x$ with the property that $-\log x$ is a 2-approximation for the Metric $k$-Center instance. To achieve this, we may use the well-known algorithms of Gonzalez [18] or Hochbaum and Shmoys [20]. We impose an arbitrary order on the vertices in $S$, i.e., we label them $s_{1},s_{2},\ldots,s_{2k}$. We then add the $12k-$ 1 edges $s_{1}s_{2},s_{1}s_{3},\ldots,s_{1}s_{2k}$. We claim that the resulting information graph $G^{\prime}$ has broadcast at least $x^{2}\alpha^{2}$. To see this, consider two arbitrary vertices $u$ and $v$. There exist (possibly non-distinct) centers $s_{u},s_{v}\in V$ such that $\mathcal{P}_{G^{\prime}}({u},{s_{u}}),\mathcal{P}_{G^{\prime}}({v},{s_{v}})\geq x$. Otherwise, there are no centers within $-\log x$ of $u$ and $v$ in the implied metric, contradicting our definition of $x$. Due to our edge additions, $\mathcal{P}_{G^{\prime}}({s_{u}},{s_{1}}),\mathcal{P}_{G^{\prime}}({s_{v}},{s_{1}})\geq\alpha$. Then we have
$\mathcal{P}_{G^{\prime}}({u},{v})\geq\mathcal{P}_{G^{\prime}}({u},{s_{u}})\cdot\mathcal{P}_{G^{\prime}}({s_{u}},{s_{1}})\cdot\mathcal{P}_{G^{\prime}}({s_{1}},{s_{v}})\cdot\mathcal{P}_{G^{\prime}}({s_{v}},{v})\geq x^{2}\alpha^{2}.$
The result follows from an application of Lemma 2, giving
$\beta({G^{\prime}})\geq x^{2}\alpha^{2}=\frac{(x^{1/2}(1+2k\alpha))^{4}\alpha^{2}}{(1+2k\alpha)^{4}}\geq\frac{(\beta^{*})^{4}\alpha^{2}}{(1+2k\alpha)^{4}}.$
3.2 A Single-Criteria Approach
We now give a single-criteria approximation algorithm for Broadcast Improvement, i.e., one that uses exactly $k$ edges. Our algorithm builds upon the $4$-approximation for Diameter Reduction [^3] Frati et al [17] study a generalization in which edges have positive integral weights and edge-additions have non-uniform cost. They call this problem Bounded Cost Minimum Diameter Edge Addition. given by Frati et al. [17]. There are several challenges to overcome since our distance metric does not provide some of the guarantees of the shortest path metric which are crucial in their approach. Hence, we need new structural insights to achieve the result. The first of these insights is formalized in the following technical lemma, which will also play a key role in Section 5. In the following, we recall the notation defined in Section 2.
Lemma 3.
Consider an information graph $G=(V,E,\alpha)$, three vertices $i,j,u\in V$, and some $0\leq\beta\leq 1$. Let $P_{u}$ be a set of paths from $i$ to $j$, all of which use vertex $u$, and let $P_{0}$ be another set of paths from $i$ to $j$ such that $P_{u}\cap P_{0}=\emptyset$ and $\Pr[P_{u}\cup P_{0}]\geq\beta$. Suppose that $\Pr[P_{0}]\leq\beta/2$. Then $\Pr[P_{u}[i,u]]\cdot\Pr[P_{u}[u,j]]\geq\beta/2$.
To understand intuitively why Lemma 3 is non-trivial, observe that the edges used by paths in $P_{u}[i,u]$ are not necessarily disjoint from those used by paths in $P_{u}[u,j]$. In this case, the existence of a path in the former segment is positively correlated with the existence of a path in the latter segment. Hence, it is not straightforward to claim that we may lower bound the product $\Pr[P_{u}[i,u]]\cdot\Pr[P_{u}[u,j]]$ by the contribution of $P_{u}$. Indeed, we suggest that completely characterizing the conditions under which this bound holds is a research direction which may contribute greatly to our understanding of information graphs. Lemma 3 sidesteps the issue by providing a lower bound in the special case where the existence of an $i,j$ path which is not in $P_{u}$ is unlikely. It turns out that this lemma is strong enough for us to obtain the algorithmic results we desire, both here and in Section 5.
Proof.
Let $\mathcal{E}_{ij}$ be the event that at least one path from $P_{u}\cup P_{0}$ exists in a graph sampled from $G$. Let $\mathcal{E}_{u}$ (resp., $\mathcal{E}_{0}$) be the event that at least one path from $P_{u}$ (resp., $P_{0}$) exists in a sampled graph. By assumption, $\Pr[\mathcal{E}_{ij}]=\Pr[\mathcal{E}_{u}\cup\mathcal{E}_{0}]\geq\beta$, and thus $\Pr[\mathcal{E}_{0}]+\Pr[\mathcal{E}_{u}\setminus\mathcal{E}_{0}]\geq\beta$. Because $\Pr[\mathcal{E}_{0}]\leq\beta/2$ (by hypothesis), we have $\Pr[\mathcal{E}_{u}\setminus\mathcal{E}_{0}]\geq\beta/2$. The main step is now to show that this implies $\Pr[P_{u}[i,u]]\cdot\Pr[P_{u}[u,j]]\geq\beta/2$.
Define $\mathcal{E}_{i}$ (resp., $\mathcal{E}_{j}$) to be the event that one of the paths from $P_{u}[i,u]$ (resp., $P_{u}[u,j]$) exists in a graph sampled from $G$. As such, the events $\mathcal{E}_{i}$ and $\mathcal{E}_{j}$ are positively correlated, i.e., $\Pr[\mathcal{E}_{i}|\mathcal{E}_{j}]\geq\Pr[\mathcal{E}_{i}]$. This implies that $\Pr[\mathcal{E}_{i}\cap\mathcal{E}_{j}]\geq\Pr[\mathcal{E}_{i}]\Pr[\mathcal{E}_{j}]$. But unfortunately, the inequality goes in the reverse direction of what we would like, as our goal is to lower-bound the product $\Pr[\mathcal{E}_{i}]\Pr[\mathcal{E}_{j}]$. To overcome this, the key is to observe that paths in $P_{u}$ are all simple (they do not contain any repeated vertices). Thus, if a path $p\in P_{u}$ exists in the sampled graph, we can find two vertex-disjoint (not considering $u$) paths $p_{i}\in P_{u}[i,u]$ and $p_{j}\in P_{u}[u,j]$ that are also in the sampled graph. Thus, conditioned on $p_{i}$, because of the disjointness requirement, we see that the existence of $p_{j}$ (and hence $p$) is actually negatively correlated. We show how to make this idea more formal below.
Let $\mathcal{E}_{ij}^{*}$ be the event that in a sampled graph, there exists $p_{i}\in P_{u}[i,u]$ and $p_{j}\in P_{u}[u,j]$ that both occur, and moreover that $p_{i},p_{j}$ do not share any vertices (other than $u$). Note that $\Pr[\mathcal{E}_{u}\setminus\mathcal{E}_{0}]=\Pr[\mathcal{E}_{u}\cap\bar{\mathcal{E}_{0}}]\leq\Pr[\mathcal{E}_{ij}^{*}\cap\bar{\mathcal{E}_{0}}]$, by definition. To complete the proof, we will show that $\Pr[\mathcal{E}_{ij}^{*}\cap\bar{\mathcal{E}_{0}}]\leq\Pr[\mathcal{E}_{i}]\Pr[\mathcal{E}_{j}]$.
At a high level, our proof strategy relies on a partition of the sample space, where the experiment in question is the sampling of $G$. In particular, we consider all subsets $O$ of $P_{u}[u,j]$, and we partition the sample space according to which subset exists in an outcome, i.e., in a graph sampled from $G$. For a particular set $O\subseteq P_{u}[u,j]$, we write $\textsf{out}_{O}$ for the event that every path in $O$ exists and no path in $P_{u}[u,j]\setminus O$ exists, meaning that for each path $p\in P_{u}[u,j]\setminus O$, at least one edge used by $p$ does not exist. Observe that the events $\textsf{out}_{O}$ partition the sample space, as desired. The proof proceeds by developing upper bounds for $(\mathcal{E}_{ij}^{*}\cap\bar{\mathcal{E}_{0}})$ conditioned on each event $\textsf{out}_{O}$, and then using the law of total probability to sum across the partition. The first step is formalized in the following claim.
Claim.
Let $O$ be any subset of $P_{u}[u,j]$, and let $\textsf{out}_{O}$ be the event defined above. Then $\Pr[(\mathcal{E}_{ij}^{*}\cap\bar{\mathcal{E}_{0}})|\textsf{out}_{O}]\leq\Pr[\mathcal{E}_{i}]$.
Proof of claim..
Let $p$ be any path from $i$ to $u$ in $P_{u}[i,u]$. We say that $p$ is uninvolved in the set $O$ if it is vertex-disjoint (not considering $u$) from every path in $O$. Let $P_{u}^{\prime}[i,u]\subseteq P_{u}[i,u]$ be the subset of paths which are uninvolved in $O$. Let $\mathcal{E}_{i}^{\prime}$ be the event that in a graph sampled from $G$ at least one path from $P_{u}^{\prime}[i,u]$ exists. Observe that, from the definitions, $\Pr[\mathcal{E}_{ij}^{*}\cap\bar{\mathcal{E}_{0}}\cap\textsf{out}_{O}]\leq\Pr[\mathcal{E}_{i}^{\prime}\cap\textsf{out}_{O}]$. Using the definition of conditional probability, we now have that $\Pr[(\mathcal{E}_{ij}^{*}\cap\bar{\mathcal{E}_{0}})|\textsf{out}_{O}]\leq\Pr[\mathcal{E}_{i}^{\prime}|\textsf{out}_{O}]$. Moreover, $\Pr[\mathcal{E}_{i}^{\prime}]\leq\Pr[\mathcal{E}_{i}]$. We complete the proof by showing that $\Pr[\mathcal{E}_{i}^{\prime}|\textsf{out}_{O}]\leq\Pr[\mathcal{E}_{i}^{\prime}]$.
Using Bayes’ Theorem, we write
$\Pr[\mathcal{E}_{i}^{\prime}|\textsf{out}_{O}]=\frac{\Pr[\textsf{out}_{O}|\mathcal{E}_{i}^{\prime}]\cdot\Pr[\mathcal{E}_{i}^{\prime}]}{\Pr[\textsf{out}_{O}]}.$
Hence, it suffices to show that $\Pr[\textsf{out}_{O}|\mathcal{E}_{i}^{\prime}]\leq\Pr[\textsf{out}_{O}]$. Intuitively, we are arguing that the event $\textsf{out}_{O}$ is not positively correlated with the existence of a path from $P_{u}[i,u]$ which is uninvolved in $O$. Formally, let $\mathcal{E}_{1}$ be the event that all paths in $O$ exist in a sampled graph, and let $\mathcal{E}_{2}$ be the event that at least one path from $P_{u}[u,j]\setminus O$ exists. Observe that $(\mathcal{E}_{1}\cap\bar{\mathcal{E}_{2}})=\textsf{out}_{O}$, and that
$\Pr[\bar{\mathcal{E}_{1}}]+\Pr[\textsf{out}_{O}]+\Pr[\mathcal{E}_{1}\cap\mathcal{E}_{2}]=1,$
and
$\Pr[\bar{\mathcal{E}_{1}}|\mathcal{E}_{i}^{\prime}]+\Pr[\textsf{out}_{O}|\mathcal{E}_{i}^{\prime}]+\Pr[(\mathcal{E}_{1}\cap\mathcal{E}_{2})|\mathcal{E}_{i}^{\prime}]=1.$
Because the edges used by paths in $P_{u}^{\prime}[i,u]$ are disjoint with those used by paths in $O$, $\mathcal{E}_{i}^{\prime}$ and $\mathcal{E}_{1}$ are independent events. That is, $\Pr[\bar{\mathcal{E}_{1}}|\mathcal{E}_{i}^{\prime}]=\Pr[\bar{\mathcal{E}_{1}}]$. To compare $\Pr[\mathcal{E}_{1}\cap\mathcal{E}_{2}]$ and $\Pr[(\mathcal{E}_{1}\cap\mathcal{E}_{2})|\mathcal{E}_{i}^{\prime}]$, we rearrange
$\Pr[(\mathcal{E}_{1}\cap\mathcal{E}_{2})|\mathcal{E}_{i}^{\prime}]=\frac{\Pr[\mathcal{E}_{1}\cap\mathcal{E}_{2}\cap\mathcal{E}_{i}^{\prime}]}{\Pr[\mathcal{E}_{i}^{\prime}]}=\Pr[\mathcal{E}_{1}|\mathcal{E}_{i}^{\prime}]\cdot\Pr[\mathcal{E}_{2}|(\mathcal{E}_{1}\cap\mathcal{E}_{i}^{\prime})].$
We have already observed that $\Pr[\mathcal{E}_{1}|\mathcal{E}_{i}^{\prime}]=\Pr[\mathcal{E}_{1}]$. Now, observe that $\Pr[\mathcal{E}_{2}|(\mathcal{E}_{1}\cap\mathcal{E}_{i}^{\prime})]\geq\Pr[\mathcal{E}_{2}|\mathcal{E}_{1}]$. In other words, the additional knowledge that at least one path from $P_{u}^{\prime}[i,u]$ exists does not decrease the likelihood of a path from $P_{u}[u,j]\setminus O$ existing. Indeed, if the edges used by paths in $P_{u}^{\prime}[i,u]$ overlap with the edges used by paths in $P_{u}[u,j]\setminus O$, then the correlation is actually positive.
We conclude that $\Pr[\bar{\mathcal{E}_{1}}|\mathcal{E}_{i}^{\prime}]=\Pr[\bar{\mathcal{E}_{1}}]$ and $\Pr[(\mathcal{E}_{1}\cap\mathcal{E}_{2})|\mathcal{E}_{i}^{\prime}]\geq\Pr[\mathcal{E}_{1}\cap\mathcal{E}_{2}]$. Consequently, $\Pr[\textsf{out}_{O}|\mathcal{E}_{i}^{\prime}]\leq\Pr[\textsf{out}_{O}]$, as desired. This completes the proof. ∎
With the claim proven, we complete the proof by using the law of total probability. That is, because the events $\textsf{out}_{O}$ partition the sample space, we have
$\Pr[\mathcal{E}_{ij}^{*}\cap\bar{\mathcal{E}_{0}}]=\sum_{O\subseteq P_{u}[u,j]}\Pr[(\mathcal{E}_{ij}^{*}\cap\bar{\mathcal{E}_{0}})|\textsf{out}_{O}]\Pr[\textsf{out}_{O}].$
For the set $O=\emptyset$, we note that the conditional probability $\Pr[(\mathcal{E}_{ij}^{*}\cap\bar{\mathcal{E}_{0}})|\textsf{out}_{O}]=0$. For all other $O$, we use the claim to write,
$\Pr[\mathcal{E}_{ij}^{*}\cap\bar{\mathcal{E}_{0}}]\leq\Pr[\mathcal{E}_{i}]\cdot\sum_{P_{u}[u,j]\supseteq O\neq\emptyset}\Pr[\textsf{out}_{O}]=\Pr[\mathcal{E}_{i}]\Pr[\mathcal{E}_{j}],$
as desired. ∎
We now describe our algorithm. Let $E^{*}$ be the optimal set of edges to add to $G$ for Broadcast Improvement, $G^{*}=(V,E\cup E^{*})$, $\beta^{*}$ be the broadcast of $G^{*}$, and $D^{*}=-\log\beta^{*}{}$. We use the algorithms of Gonzalez [18] or Hochbaum and Shmoys [20] to obtain a 2-approximation for the Metric $k$-Center instance on the implied metric $(V,\phi)$ of $G$ with parameter $k+1$. If we let $M_{k+1}$ be the optimal radius of the Metric $k$-Center instance and $\mu_{k+1}$ be the value returned by the algorithm then $\mu_{k+1}\leq 2M_{k+1}$. Let $X\subseteq V$ be the set of $k+1$ vertices returned by the 2-approximation algorithm. Using the vertices of $X$ add a star in $G$ (with the center chosen arbitrarily from $X$) and let $G^{\prime}$ be this new graph. Thus, we added at most $k$ new edges to $G$ and we will show that the broadcast of $G^{\prime}$ (call it $\beta^{\prime}$ and let $D^{\prime}=-\log\beta^{\prime}$) is at least $\frac{\beta^{*}{}^{4}\alpha^{2}}{16^{k+1}}$.
For every vertex $s$ in $G^{\prime}$ (and also in $G$), at least one of the centers from $X$ is within distance $\mu_{k+1}$ in the implied metric of $G$, and therefore also in the implied metric of $G^{\prime}$. Thus, in the implied metric of $G^{\prime}$, starting from any vertex $s$, by reaching the nearest center in distance at most $\mu_{k+1}$, then taking one edge to reach the root of the star we added, taking one more edge to the center closest to the destination vertex $t$, and then traveling a distance of at most $\mu_{k+1}$ again to reach $t$ – in at most $2\mu_{k+1}-2\log\alpha$ distance, one can reach any vertex in the graph starting from any vertex. Thus,
$D^{*}\leq D^{\prime}\leq 2\mu_{k+1}-2\log\alpha\leq 2(2M_{k+1})-2\log\alpha\leq 4M_{k+1}-2\log\alpha$ (1)
We will now show that $M_{k+1}$ is upper-bounded by $D^{*}+k+1$. Combined with Eq. 1 this allows us to bound the approximation factor.
To show that $M_{k+1}\leq D^{*}+k+1$, we will show that $G$ is a feasible instance of Metric $k$-Center for parameter value $k+1$ and distance to center at most $D^{*}+k+1$. We pick the centers in the following way: we first pick an arbitrary vertex $x$ as a center. Consider the set of at most $2k$ vertices that are the end points of $E^{*}$. For every new edge $(i,j)$ in $G^{*}$, if $\mathcal{P}_{G^{*}\setminus(i,j)}({x},{i})\geq\mathcal{P}_{G^{*}\setminus(i,j)}({x},{j})$ then we pick $j$ as a center. Otherwise, we pick $i$ as a center. In other words, we pick as a center whichever of $i$ and $j$ is farther from $x$ in the implied metric of $G^{*}\setminus(i,j)$. In this way, we get a set of at most $k+1$ centers. Call this set $\hat{X}$. We now prove that $\hat{X}$ is a feasible solution.
Lemma 4.
For every vertex $w\in V$ at least one of the vertices in $\hat{X}$ is within distance $D^{*}+k+1$ of $w$ in the implied metric of $G$.
Proof.
If $w$ is a center, the claim is trivially true. So, assume $w$ is not a center. Let $x$ be an arbitrary vertex, and let $P_{G^{*}}(w,x)$ be the set of all simple paths from $w$ to $x$ in $G^{*}$. Call the edges in $E^{*}$ new edges. Consider the partition of $P_{G^{*}}(w,x)$ into $12k+$ 1 equivalence classes based on the first new (directed) edge they encounter going from $w$ to $x$. Let $C_{vu}$ represent the class of simple $(w,x)$ paths such that $(u,v)$ is the first new edge they encounter and $v$ appears before $u$ on these paths. Define $C_{uv}$ similarly. Suppose in the new edge $(u,v)$, $v$ is the center. Then, we will say that $C_{vu}$ and $C_{uv}$ are the center-first and the center-second classes, respectively. Moreover, since $v$ is a center, $\mathcal{P}_{G^{*}\setminus(u,v)}({x},{v})\leq\mathcal{P}_{G^{*}\setminus(u,v)}({x},{u})$.
If the contribution of the center-first class $C_{vu}$ is at least $\frac{\beta^{*}{}}{2^{k+1}}$ then $\mathcal{P}_{G}({w},{v})\geq\frac{\beta^{*}{}}{2^{k+1}}$ and the claim is true. So, assume $\Pr[C_{vu}]\leq\frac{\beta^{*}{}}{2^{k+1}}$. Then,
$\Pr[P_{G^{*}}(w,x)\setminus(C_{uv}\cup C_{vu})]+\Pr[C_{uv}]\geq\beta^{*}{}-\frac{\beta^{*}}{2^{k+1}}$
We now claim that the first summand accounts for most of this probability mass.
Claim.
$\Pr[P_{G^{*}}(w,x)\setminus(C_{uv}\cup C_{vu})]\geq\frac{1}{2}(\beta^{*}{}-\frac{\beta^{*}}{2^{k+1}})$.
Proof of claim..
Suppose otherwise. Then Lemma 3 gives us that
$\Pr[C_{uv}[w,u]]\cdot\Pr[C_{uv}[u,x]]=\Pr[C_{uv}[w,u]]\cdot\alpha\cdot\Pr[C_{uv}[v,x]]\geq\frac{1}{2}(\beta^{*}{}-\frac{\beta^{*}}{2^{k+1}}).$
By noting that $\mathcal{P}_{G^{*}\setminus(u,v)}({w},{u})\geq\Pr[C_{uv}[w,u]]$ and $\mathcal{P}_{G^{*}\setminus(u,v)}({v},{x})\geq\Pr[C_{uv}[v,x]]$, we observe that
$\mathcal{P}_{G^{*}\setminus(u,v)}({w},{u})\cdot\mathcal{P}_{G^{*}\setminus(u,v)}({v},{x})\geq\frac{1}{2}(\beta^{*}{}-\frac{\beta^{*}}{2^{k+1}}).$
We can now manipulate definitions to obtain a contradiction.
$&\Pr[P_{G^{*}}(w,x)\setminus(C_{uv}\cup C_{vu})]\\ &\geq\Pr[P_{G^{*}\setminus(u,v)}(w,x)]\\ &=\mathcal{P}_{G^{*}\setminus(u,v)}({w},{x})\\ &\geq\mathcal{P}_{G^{*}\setminus(u,v)}({w},{u})\cdot\mathcal{P}_{G^{*}\setminus(u,v)}({u},{x})\\ &\geq\mathcal{P}_{G^{*}\setminus(u,v)}({w},{u})\cdot\mathcal{P}_{G^{*}\setminus(u,v)}({v},{x})\\ &\geq(1/2)(\beta^{*}{}-\beta^{*}/2^{k+1})\\ \end{aligned}\begin{aligned} &\\ \quad&(\text{\hfill since }P_{G^{*}\setminus(u,v)}(w,x)\subseteq P_{G^{*}}(w,x)\setminus(C_{uv}\cup C_{vu}))\\ &(\text{by definition of proximity})\\ &(\text{by triangle inequality of the implied metric})\\ &(\text{since }\mathcal{P}_{G^{*}\setminus(u,v)}({v},{x})\leq\mathcal{P}_{G^{*}\setminus(u,v)}({u},{x}))\\ &(\text{by our argument above}),\\ \end{aligned}$
but we began this proof by supposing that $\Pr[P_{G^{*}}(w,x)\setminus(C_{uv}\cup C_{vu})]<\frac{1}{2}(\beta^{*}{}-\frac{\beta^{*}}{2^{k+1}})$.
We repeat the same analysis by looking at the now remaining $12k-$ 1 classes. If $(u^{\prime},v^{\prime})$ is another new edge and the center-first class of $(u^{\prime},v^{\prime})$ has contribution at most $\frac{\beta^{*}}{2^{k+1}}$ then the remaining $32k-$ 3 classes must have contribution at least
$\frac{1}{2}(\frac{\beta^{*}{}}{2}(1-\frac{1}{2^{k}})-\frac{\beta^{*}{}}{2^{k+1}})=\frac{\beta^{*}{}}{4}(1-\frac{1}{2^{k}}-\frac{1}{2^{k}})=\frac{\beta^{*}{}}{4}(1-\frac{1}{2^{k-1}}).$
Thus, if after $t$ rounds the contribution of the remaining $12(k-t)+$ 1 classes is $\frac{\beta^{*}{}}{2^{t}}(1-\frac{1}{2^{k-t+1}})$ then after 1 more round the contribution of the remaining $12(k-t)-$ 1 classes will be at least $\frac{\beta^{*}{}}{2^{t+1}}(1-\frac{1}{2^{k-(t+1)+1}})$.
Therefore, if the center-first classes of none of the $k$ new edges have contribution at least $\frac{\beta^{*}{}}{2^{k+1}}$ in $G^{*}$ then the “empty class” (corresponding to paths that have no new edge) must have contribution $\geq\frac{\beta^{*}{}}{2^{k}}(1-\frac{1}{2})\geq\frac{\beta^{*}{}}{2^{k+1}}$ in $G^{*}$ (and hence $G$). This completes the proof. ∎
Putting it all together, we get the main result.
Theorem 2.
There exists a polynomial-time algorithm which produces an information graph with broadcast at least $\frac{(\beta^{*})^{4}\alpha^{2}}{16^{k+1}}$ by adding at most $k$ edges.
3.3 Harnessing Bounded Doubling Dimension
We conclude this section by noting that an improved Metric $k$-Center approximation immediately improves the guarantees given by Theorems 1 and 2. Unfortunately, it is NP-hard to approximate Metric $k$-Center below factor 2 [21]. In the following, we describe how to improve the approximation guarantee if one is willing to accept runtime which is fixed-parameter tractable in the combination of $k$ and the doubling dimension of the implied metric.
Definition 1.
The doubling dimension of a metric space $X$ is the smallest number $d$ such that, for every $r>0$, any ball in $X$ of radius $r$ is contained in the union of at most $2^{d}$ balls of radius $r/2$.
Here, we make use of an efficient parameterized approximation scheme for Metric $k$-Center due to Feldmann and Marx [15].
Theorem 3 ([15]).
Given a metric space of doubling dimension $d$ and $\varepsilon>0$, a $(1+\varepsilon)$-approximation for Metric $k$-Center can be computed in $(k^{k}/\varepsilon^{O(kd)})\cdot n^{O(1)}$ time.
Using Lemma 2 and the same edge addition strategies as in Theorems 1 and 2, we can obtain a parameterized approximation for Broadcast Improvement.
Theorem 4.
For any $\alpha,\varepsilon>0$, there exist algorithms producing information graphs with broadcast at least (i) $\frac{(\beta^{*})^{2+\varepsilon}\alpha^{2}}{(1+2k\alpha)^{2+\varepsilon}}$ by adding at most $12k-$ 1 edges or (ii) $\frac{(\beta^{*})^{2+\varepsilon}\alpha^{2}}{16^{k+1}}$ by adding at most $k$ edges, in $(4^{k}k^{2k}2^{O(kd)}/\varepsilon^{O(kd)})\cdot n^{O(1)}$ time, where $d$ is the doubling dimension of the implied metric $(V,\phi)$.
It is natural to ask how large the doubling dimensions of information graphs’ implied metrics tend to be. A trivial upper bound is $\log n$. In general, this bound is tight up to factor 2, both in very dense and very sparse information graphs. Consider a clique, and let $x$ be the pairwise proximity of any two nodes in this clique. In the implied metric, any ball of radius $-\log x$ contains every vertex. However, any ball of radius $-\log\sqrt{x}$ contains only a single point, its center. Thus, the doubling dimension is $\log n$. On the other hand, consider a complete binary tree of depth $d$. In the implied metric, a ball of radius $r=-\log\alpha^{d}$ around the root will contain all vertices. Now, let $u$ and $v$ be any two vertices at depth $\frac{d}{2}$, and let $u^{\prime}$ and $v^{\prime}$ be any two leaf vertices in the subtrees rooted at $u$ and $v$, respectively. It is easy to check that no ball of radius $\frac{r}{2}=-\log\alpha^{d/2}$ can contain both $u^{\prime}$ and $v^{\prime}$. Consequently, since there are $2^{d/2}$ vertices at depth $\frac{d}{2}$, the doubling dimension is at least $\frac{1}{2}\log n$.
Identifying structural and/or application domain restrictions which lead to low doubling dimensions in the implied metrics of information graphs is an interesting direction for future research. When considering shortest path metrics, previous work has argued that many graphs (e.g., transportation networks) can be assumed to be embeddable in spaces of low doubling dimension [15]. It remains to be seen for what classes, if any, of information graphs a similar assumption might be reasonable.
4 A Linear Approximation Based on Constant Witnesses
In this section we introduce $(c,b)$-witnessing solutions, and use them to obtain linear approximations for Broadcast Improvement using poly$(k)\cdot\log n$ edges.
Definition 2.
Given an information graph $G=(V,E,\alpha)$, two vertices $u,v\in V$, a non-negative integer $c$, and $0\leq b\leq 1$, a $(u,v,b)$-witness of size $c$ is a set $W_{uv}\subseteq V^{2}$ of size $c$ with the property that $\mathcal{P}_{G+W_{uv}}({u},{v})\geq b$. If $S\subseteq V^{2}\setminus E$ contains as a subset a $(u,v,b)$-witness of size at most $c$ for every pair of vertices $u,v\in V$, we say $S$ is a $(c,b)$-witnessing solution to Broadcast Improvement on $G$.
The main idea of our algorithm is to show that there always exists a $(c,f(\beta^{*}))$-witnessing solution of bounded size for some constant $c$ and function $f$. This allows us to use Hitting Set [^4] Given a set family $\mathcal{F}$ over a universe $\mathcal{U}$, choose a set $X\subseteq\mathcal{U}$ of minimum size such that $S\cap X\neq\emptyset$ for all $S\in\mathcal{F}$. This problem is equivalent to Set Cover.-like techniques to solve Broadcast Improvement. The key lemma is the following:
Lemma 5.
For any instance $(G=(V,E,\alpha),k)$ of Broadcast Improvement with optimum broadcast $\beta^{*}$, there exist both (i) a $(3,\frac{4\beta^{*}}{12k^{4}+3k^{2}})$-witnessing solution of size at most $67k-$ 6, and (ii) a $(2,\frac{\beta^{*}\alpha}{12k^{2}+3})$-witnessing solution of size at most $\binom{2k}{2}$.
$u$$l$$t$$v$$w$$u$$l$$t$$v$$w$
Figure 2: Two figures illustrating the construction of $Q$ in the proof of part (i) of Lemma 5. Both figures depict $\mathcal{C}^{\prime}$, which is a set of paths from $u$ to $v$ in $G+S^{*}$. On the left, $\mathcal{C}^{\prime}=\mathcal{C}_{3}$. On the right, $\mathcal{C}^{\prime}=\mathcal{C}_{1}$ (the case in which $\mathcal{C}^{\prime}=\mathcal{C}_{2}$ is symmetric). Red edges denote the fan-out and fan-in edges of $\mathcal{C}^{\prime}$. Blue edges are the middle segment of $Q$.
Proof.
We begin with the proof of part (i). Let $S^{*}$ be an optimum solution producing broadcast $\beta^{*}$ in the information graph $G^{*}=G+S^{*}$. We begin by imposing an arbitrary ordering on the endpoints of the edges in $S^{*}$, $s_{1},s_{2},\ldots,s_{m}$, where $m\leq 2k$. We construct a new solution $S$, where
$S=S^{*}\cup\{s_{1}s_{2},s_{1}s_{3},\ldots s_{1}s_{m}\}\cup\{s_{2}s_{3},s_{2}s_{4},\ldots s_{2}s_{m}\}\cup\{s_{3}s_{4},s_{3}s_{5},\ldots s_{3}s_{m}\}$
Intuitively, we have chosen three [^5] Note that the edges of $S^{*}$ always involve at least three distinct endpoints, unless $k=1$. In this latter case, Broadcast Improvement is polynomial-time solvable. distinct endpoints of edges in $S^{*}$, and formed stars with these endpoints as the centers and all other endpoints as the leaves. It is easy to check that $S$ has size at most $k+3m-6\leq 7k-6$. To complete the proof, we must show that $S$ contains a $(u,v,\frac{4\beta^{*}}{12k^{4}+3k^{2}})$-witness of size at most three for every pair of vertices $u,v\in V$.
Let $P_{uv}$ be the set of paths from $u$ to $v$ in $G^{*}$, and recall (see Section 2) that the contribution of $P_{uv}$ is exactly equal to $\mathcal{P}_{G^{*}}({u},{v})$, and therefore an upper bound for $\beta^{*}$. Now, let $p_{i}$ and $p_{j}$ be two paths in $P_{uv}$. Let $l_{i}$ be the leading vertex of the first edge contained in $S^{*}$ (a new edge) to appear along $p_{i}$. Also, let $t_{i}$ be the trailing vertex of the last new edge to appear along $p_{i}$. Define $l_{j}$ and $t_{j}$ similarly. We impose an equivalence relation on $P_{uv}$ by declaring that $p_{i}$ is similar to $p_{j}$ if $l_{i}=l_{j}$ and $t_{i}=t_{j}$. Note that we may reserve one equivalence class for the set of paths containing no new edges, so the equivalence relation remains well-defined. This relation partitions $P_{uv}$ into at most $14k^{2}+$ 1 equivalence classes, and the sum of the contributions of these classes is an upper bound for $\beta^{*}$. It follows that there must be at least one class of paths with contribution at least $\frac{\beta^{*}}{4k^{2}+1}$. If the class of paths containing no new edges meets this criteria, then we are done, as the empty set is a $(u,v,\frac{\beta^{*}}{4k^{2}+1})$-witness, and $\frac{\beta^{*}}{4k^{2}+1}>\frac{4\beta^{*}}{12k^{4}+3k^{2}}$ as long as $k>1$. Otherwise, choose one such equivalence class, defined by vertices $l$ and $t$, and call this class $\mathcal{C}$. At least one of $s_{1},s_{2}$, or $s_{3}$ is distinct from both $l$ and $t$. Call this vertex $w$.
We now further partition $\mathcal{C}$ into three subsets. The first, denoted $\mathcal{C}_{1}$, is the set of paths in $\mathcal{C}$ on which $w$ precedes $l$. The second, $\mathcal{C}_{2}$, is those paths on which $t$ precedes $w$. The third, $\mathcal{C}_{3}$, is all other paths in $\mathcal{C}$. The sum of the contributions of these three subsets is an upper bound for the contribution of $\mathcal{C}$. It follows that at least one has contribution at least $\frac{\beta^{*}}{3(4k^{2}+1)}$. Let $\mathcal{C}^{\prime}$ denote whichever of $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, or $\mathcal{C}_{3}$, has the largest contribution. We will now show how to replace $\mathcal{C}^{\prime}$ with a new set of paths $Q$ which uses at most three edges from $S$.
We begin by handling a special case, namely the case in which the edge $lt$ appears along at least one path in $\mathcal{C}^{\prime}$. In this case, we use the fact that $lt\in S$ to include those paths in $Q$ with no modifications. Consequently, conditioned on the existence of the edge $lt$ in a sampled graph, the existence of any path in $\mathcal{C}^{\prime}$ implies the existence of a path in $Q$. In other words, when conditioned on the existence of the edge $lt$, the contribution of $Q$ is at least as large as the contribution of $\mathcal{C}^{\prime}$. It is therefore sufficient for us to show that the contributions of $Q$ and $\mathcal{C}^{\prime}$ are not too different when conditioned on the non-existence of $lt$ in a sampled graph. Hence, we proceed with the simplifying assumption that the edge $lt$ does not appear along any path in $\mathcal{C}^{\prime}$.
We now show how to edit these paths to form $Q$. See Figure 2 for a visual aid. If $\mathcal{C}^{\prime}=\mathcal{C}_{1}$, then for each $p\in\mathcal{C}^{\prime}$ we replace $p[w,t]$ with the edge $wt$. If $\mathcal{C}^{\prime}=\mathcal{C}_{2}$, then for each $p\in\mathcal{C}^{\prime}$ we replace $p[l,w]$ with the edge $lw$. Otherwise, we replace $p[l,t]$ with the segment $(l,lw,w,wt,t)$. We observe that if $\mathcal{C}^{\prime}=\mathcal{C}_{1}$, then $\mathcal{C}^{\prime}[u,w]=Q[u,w]$ and $\mathcal{C}^{\prime}[t,v]=Q[t,v]$. Similarly, if $\mathcal{C}^{\prime}=\mathcal{C}_{2}$ then $\mathcal{C}^{\prime}[u,l]=Q[u,l]$ and $\mathcal{C}^{\prime}[w,v]=Q[w,v]$, and if $\mathcal{C}^{\prime}=\mathcal{C}_{3}$ then $\mathcal{C}^{\prime}[u,l]=Q[u,l]$ and $\mathcal{C}^{\prime}[t,v]=Q[t,v]$. We call the segment of $\mathcal{C}^{\prime}$ on which $Q$ differs, namely either $\mathcal{C}^{\prime}[w,t],\mathcal{C}^{\prime}[l,w]$, or $\mathcal{C}^{\prime}[l,t]$ the middle segment of $\mathcal{C}^{\prime}$, denoted $\mathcal{C}^{\prime}_{m}$, and we call the other two segments the beginning and ending segments, written $\mathcal{C}^{\prime}_{b}$ and $\mathcal{C}^{\prime}_{e}$, respectively. We define $Q_{b},Q_{m}$, and $Q_{e}$ similarly. Moreover, we note that whatever the value of $\mathcal{C}^{\prime}\in\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$, $Q_{b}=\mathcal{C}^{\prime}_{b}$ and $Q_{e}=\mathcal{C}^{\prime}_{e}$.
Thus far, we have identified a set of paths $\mathcal{C}^{\prime}$ from $u$ to $v$ in $G+S^{*}$ with contribution at least $\frac{\beta^{*}}{12k^{2}+3}$, and we have used $\mathcal{C}^{\prime}$ to construct a new set $Q$ of paths which use (in total) at most three edges from $S$. To complete the proof, it is sufficient to show that $\Pr[Q]\geq\frac{4}{k^{2}}\Pr[\mathcal{C}^{\prime}]$ (recalling the notation defined in Section 2 for the contribution of a set of paths). Intuitively, we accomplish this first by arguing that since $Q$ and $\mathcal{C}^{\prime}$ have identical beginning and ending segments, it is sufficient to compare their middle segments, and second by performing that comparison. However, the potential positive correlation between paths in different segments of $\mathcal{C}^{\prime}$ necessitates a slightly more technical argument.
Let $(p_{1},p_{2})$ be a pair of paths from the beginning and ending segments of $\mathcal{C}^{\prime}$, i.e., $p_{1}\in\mathcal{C}^{\prime}_{b}=Q_{b}$ and $p_{2}\in\mathcal{C}^{\prime}_{e}=Q_{e}$. We say that $(p_{1},p_{2})$ is a nice path pair if $p_{1}$ and $p_{2}$ are vertex-disjoint, and that $(p_{1},p_{2})$ exists in a sampled graph if both paths exist. Let $\mathcal{E}_{1}$ be the event that a nice path pair exists in a sampled graph666 Technically, $\mathcal{E}_{1}$ is an event in two sample spaces, i.e., the spaces defined by sampling from $G+S^{*}$ and $G+S$. However, since edges are sampled independently and the edges relevant to $\mathcal{E}_{1}$ exist in both graphs, the event remains well-defined. . Note that by construction, the vertex $w$ does not appear on any path in either $Q_{b}$ or $Q_{e}$. Then the edges of the paths in $Q_{m}$ are disjoint from the edges of paths in $Q_{b}$ and $Q_{e}$. Noting that edges are sampled independently, we now have that $\Pr[Q]=\Pr[Q_{m}]\cdot\Pr[\mathcal{E}_{1}]$. Moreover, because $Q_{m}$ consists of a single path on at most two edges, i.e., either the edge $wt$, the edge $lw$, or the path $(l,lw,w,wt,t)$, we may write $\Pr[Q_{m}]\geq\alpha^{2}$, and conclude that $\Pr[Q]\geq\alpha^{2}\Pr[\mathcal{E}_{1}]$.
We now upper bound $\Pr[\mathcal{C}^{\prime}]$. Unfortunately, the existence (in a sampled graph) of a path in $\mathcal{C}^{\prime}_{m}$ may be positively correlated to the existence of a nice path pair, so it is not straightforward to claim that $\Pr[\mathcal{C}^{\prime}]\leq\Pr[\mathcal{C}^{\prime}_{m}]\cdot\Pr[\mathcal{E}_{1}]$. Instead, we will leverage the structure imposed by our definition of $\mathcal{C}^{\prime}$ to obtain another upper bound which is sufficient for our purposes. We call those new edges (edges in $S^{*}$) which are incident to $l$ and used by at least one path in $\mathcal{C}^{\prime}$ the fan-out edges of $\mathcal{C}^{\prime}$. Similarly, we call those new edges which are incident to $t$ and used by at least one path in $\mathcal{C}^{\prime}$ the fan-in edges of $\mathcal{C}^{\prime}$. Let $\mathcal{E}_{out}$ (respectively, $\mathcal{E}_{in}$) be the event that at least one fan-out (respectively, fan-in) edge exists in a sampled graph. Observe that $\Pr[\mathcal{C}^{\prime}]\leq\Pr[\mathcal{E}_{1}\cap\mathcal{E}_{out}\cap\mathcal{E}_{in}]$. Using the fact that, by construction, no new edges appear on any paths in $\mathcal{C}^{\prime}_{b}$ or $\mathcal{C}^{\prime}_{e}$, we have that $\mathcal{E}_{1}$ and $(\mathcal{E}_{out}\cap\mathcal{E}_{in})$ are independent. Then $\Pr[\mathcal{C}^{\prime}]\leq\Pr[\mathcal{E}_{1}]\cdot\Pr[\mathcal{E}_{out}\cap\mathcal{E}_{in}]$.
We now need only to upper bound $\Pr[\mathcal{E}_{out}\cap\mathcal{E}_{in}]$. Observe that each pair of edges, one being a fan-out edge and the other being a fan-in edge, exists in a sampled graph with probability $\alpha^{2}$. Recall, according to our prior argument, that the edge $lt$ does not appear on any path in $\mathcal{C}^{\prime}$, so the fan-out and fan-in edges are disjoint sets. Furthermore, their union has size at most $k$. We use these facts to obtain the following bound:
$\Pr[\mathcal{E}_{out}\cap\mathcal{E}_{in}]\leq\alpha^{2}\cdot\max_{i\in[k]}\{i\cdot(k-i)\}\leq\frac{\alpha^{2}k^{2}}{4}.$
Putting the whole proof together, we see that
$\mathcal{P}_{G+W_{uv}}({u},{v})\geq\Pr[Q]\geq\alpha^{2}\Pr[\mathcal{E}_{1}]=\alpha^{2}\Pr[\mathcal{E}_{1}]\cdot\frac{\Pr[\mathcal{E}_{out}\cap\mathcal{E}_{in}]}{\Pr[\mathcal{E}_{out}\cap\mathcal{E}_{in}]}\geq\frac{4}{k^{2}}\Pr[\mathcal{C}^{\prime}]\geq\frac{4\beta^{*}}{12k^{4}+3k^{2}},$
where $W_{uv}$ consists of the at most three edges from $S$ appearing along paths in $Q$. Hence, we have found the desired witness.
To prove part (ii) of the lemma, we use a larger solution. Recalling that we have labeled the endpoints of the edges in $S^{*}$ $s_{1},s_{2},\ldots$ we create a new solution $S=\{s_{i}s_{j}\ |i\neq j\}$ of size at most $\binom{2k}{2}$. That is, we add an edge between each pair of endpoints of edges in $S^{*}$. The next part of the proof proceeds as before, up to the definition of $\mathcal{C}^{\prime}$. Note that this time we do not need to handle the special case concerning edge $lt$ separately. We now construct our set $Q$ by replacing the segment $\mathcal{C}^{\prime}[l,t]$ with the edge $lt$. This is the only new edge used by $Q$, so all that remains is to show that $\Pr[Q]\geq\frac{\beta^{*}\alpha}{12k^{2}+3}$. To accomplish this, we use the same definitions as before for nice path pairs and the event $\mathcal{E}_{1}$. We note that $\Pr[\mathcal{C}^{\prime}]\leq\Pr[\mathcal{E}_{1}]$ by definition, and $\Pr[Q]=\alpha\Pr[\mathcal{E}_{1}]$ because the sampling of edge $lt$ occurs independently of event $\mathcal{E}_{1}$. The claim follows.
We use Lemma 5 to form a reduction from Broadcast Improvement to Hitting Set. Additionally, we show how to repeat this reduction in a binary search fashion to find a good estimate for the value of $\beta^{*}{}$.
Theorem 5.
For any $\varepsilon,\alpha>0$, there exist polynomial-time algorithms which produce information graphs with broadcast at least (i) $\frac{4\beta^{*}}{(1+\varepsilon)(12k^{4}+3k^{2})}$ using $O(k^{3}\log n)$ edge additions, and (ii) $\frac{\beta^{*}\alpha}{(1+\varepsilon)(12k^{2}+3)}$ using $O(k^{2}\log n)$ edge additions.
Proof.
We prove part (i) of the theorem. The proof for part (ii) is conceptually identical and therefore omitted for brevity. We will begin by assuming that we already know the value of $\beta^{*}$. In this case, we reduce to Hitting Set as follows. We define $\mathcal{W}=\binom{V^{2}\setminus E}{3}\cup\binom{V^{2}\setminus E}{2}\cup\binom{V^{2}\setminus E}{1}$ as the set containing all groups of at most three potential edge additions. Note that $|\mathcal{W}|\in O(n^{6})$. The elements of $\mathcal{W}$ are the elements of our hitting set instance. Then, for each pair of vertices $u,v\in V$ with $\mathcal{P}_{G}({u},{v})<\frac{4\beta^{*}}{12k^{4}+3k^{2}}$, we add a set $\mathcal{W}_{uv}$ consisting of all $(u,v,\frac{4\beta^{*}}{12k^{4}+3k^{2}})$-witnesses of size at most three. This completes the construction. We use Lemma 5 to observe that there exists a hitting set of size at most $\binom{7k-6}{3}+\binom{7k-6}{2}+\binom{7k-6}{1}\in O(k^{3})$.
The algorithm proceeds by using the well-known greedy $O(\log n)$-approximation for Hitting Set [24] to generate a hitting set of size $O(k^{3}\log n)$. We return the union of all the witnesses contained in this hitting set. By construction, this set of edge additions contains as a subset a $(u,v,\frac{4\beta^{*}}{12k^{4}+3k^{2}})$-witness for every pair $u,v$ of vertices, and because every member of our hitting set contains at most three edges, our solution has size $O(k^{3}\log n)$.
It remains to show how we can estimate $\beta^{*}$. We will do this by mimicking the technique of Demaine and Zadimoghaddam [11]. In the following, let $b$ denote the precise bound on edge additions given by the algorithm in the preceding paragraph. That is, $b$ is $\binom{7k-6}{3}+\binom{7k-6}{2}+\binom{7k-6}{1}\in O(k^{3})$ multiplied by the approximation factor given by [24]. We note that $\beta^{*}\leq 1$, so $\frac{\beta^{*}}{\beta({G})}\leq\frac{1}{\beta({G})}$. Then for any $\varepsilon>0$, there exists some integer $0\leq i\leq\log_{1+\varepsilon}\frac{1}{\beta({G})}$ with the property that $\beta({G})(1+\varepsilon)^{i}\leq\beta^{*}\leq\beta({G})(1+\varepsilon)^{i+1}$. We conduct a binary search of integers in the interval $[0,\log_{1+\varepsilon}\frac{1}{\beta({G})}]$. Note that $\beta({G})\geq\alpha^{n}$, so this interval has polynomial length (for fixed $\varepsilon$ and $\alpha$) [^7] Algorithms using this binary search procedure technically have running time XP in $\alpha$, if $\alpha$ is viewed as an arbitrary parameter of the input. However, in practice we believe that $\alpha$ can be modeled as some application-specific constant. We are unaware of any motivation to study Broadcast Improvement in the setting where $\alpha$ is not bounded below by some constant greater than 00.. For each tested integer $j$, we assume that $\beta^{*}=\beta({G})(1+\varepsilon)^{j}$, and execute the algorithm described above. If the algorithm adds more than $b$ edges, then we conclude that $\beta({G})(1+\varepsilon)^{j}\geq\beta^{*}$, and therefore that $j>i$. Let $j^{*}$ be the largest integer in the interval for which our algorithm adds at most $b$ edges. Then we can conclude that $\beta^{*}\leq\beta({G})(1+\varepsilon)^{j^{*}+1}$, and in this case our algorithm adds at most $b\in O(k^{3}\log n)$ edges to produce broadcast at least
$\frac{4\beta({G})(1+\varepsilon)^{j^{*}}}{12k^{4}+3k^{2}}=\frac{4\beta({G})(1+\varepsilon)^{j^{*}+1}}{(1+\varepsilon)(12k^{4}+3k^{2})}\geq\frac{4\beta^{*}}{(1+\varepsilon)(12k^{4}+3k^{2})},$
as desired. ∎
5 Reducing Edge Additions via Submodularity
Building on the techniques of Section 4, we now show how to obtain a similar broadcast guarantee while reducing the number of edge additions to $O(k\log n)$. We begin by showing that there always exists a solution of bounded suboptimality in which the added edges form a star. This claim can be proven using techniques very similar to those of Lemma 5.
Theorem 6.
Given any instance $(G,k)$ of Broadcast Improvement, there exist $12k-$ 1 edge additions $E^{\prime}$ all incident to a shared endpoint such that $\beta({G+E^{\prime}})\geq\frac{\beta^{*}\alpha^{2}}{12k^{2}+3}$.
Proof.
The proof proceeds very similarly to that of Lemma 5, but with slightly simpler analysis. Let $S^{*}$ be an optimum solution producing broadcast $\beta^{*}$ in the information graph $G^{*}=G+S^{*}$. We begin by imposing an arbitrary ordering on the endpoints of the edges in $S^{*}$: $s_{1},s_{2},\ldots,s_{m}$, where $m\leq 2k$. We construct a new solution $S=\{s_{1}s_{2},s_{1}s_{3},\ldots,s_{1}s_{m}\}$. We call $s_{1}$ $w$, and note that it is the center of our constructed star.
The next part of the proof proceeds as in Lemma 5, up to the construction of $Q$. Note that $w$ may not be distinct from $l$ and $t$, but this is not a problem, as we no longer need to handle the special case concerning edge $lt$ separately. It remains to show that the edges from $S$ used by $Q$ are a $(u,v,\frac{\beta^{*}\alpha^{2}}{12k^{2}+3})$-witness. Since $\mathcal{C}^{\prime}$ is defined in the same way as in the proof of Lemma 5, it suffices to show that $\Pr[Q]\geq\alpha^{2}\cdot\Pr[\mathcal{C}^{\prime}]$.
Once again using the notation of the proof of Lemma 5, we observe that if $\mathcal{C}^{\prime}=\mathcal{C}_{1}$, then $\mathcal{C}^{\prime}[u,w]=Q[u,w]$ and $\mathcal{C}^{\prime}[t,v]=Q[t,v]$. Similarly, if $\mathcal{C}^{\prime}=\mathcal{C}_{2}$ then $\mathcal{C}^{\prime}[u,l]=Q[u,l]$ and $\mathcal{C}^{\prime}[w,v]=Q[w,v]$, and if $\mathcal{C}^{\prime}=\mathcal{C}_{3}$ then $\mathcal{C}^{\prime}[u,l]=Q[u,l]$ and $\mathcal{C}^{\prime}[t,v]=Q[t,v]$. We once again call the segment of $\mathcal{C}^{\prime}$ on which $Q$ differs, namely either $\mathcal{C}^{\prime}[w,t],\mathcal{C}^{\prime}[l,w]$, or $\mathcal{C}^{\prime}[l,t]$ the middle segment of $\mathcal{C}^{\prime}$, and we call the other two segments the beginning and ending segments. We apply the same names to the corresponding segments of $Q$. So, whatever the value of $\mathcal{C}^{\prime}\in\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$, $Q$ and $\mathcal{C}^{\prime}$ have identical beginning and ending segments. We define nice path pairs and the event $\mathcal{E}_{1}$ as in the proof of Lemma 5. We note that $\Pr[\mathcal{C}^{\prime}]\leq\Pr[\mathcal{E}_{1}]$, by definition. Next, we observe that, according to our construction, the edges used by the middle segment of $Q$ (namely, $wt$, $lw$, or both) are disjoint from the edges used by the paths in the beginning and ending segments of $Q$. Moreover, the middle segment of $Q$ has contribution at least $\alpha^{2}$. Thus, we have that
$\Pr[Q]\geq\alpha^{2}\Pr[\mathcal{E}_{1}]\geq\alpha^{2}\Pr[\mathcal{C}^{\prime}]\geq\frac{\beta^{*}\alpha^{2}}{12k^{2}+3},$
completing the proof.
Our idea in this section is to define an appropriate function and use its submodularity to obtain an algorithm. Unfortunately, the natural candidate —improvement in broadcast when a set of edges is added— is not a submodular function.
Observation 1.
Neither the broadcast function $\beta\colon 2^{V^{2}}\rightarrow[0,1]$ nor the logarithm of the broadcast function is submodular with respect to edge additions.
To understand 1, it is helpful to consider a small example: a sub-divided star with three leaves. That is, our graph $G$ consists of a “center” vertex $v_{0}$ connected to each of three leaves $v_{1},v_{2}$, and $v_{3}$ via disjoint paths on $\ell$ edges, where $\ell$ can be thought of as some large integer whose value depends on $\alpha$. The broadcast of $G$ is $\alpha^{2\ell}$. Now, we observe that $G+\{v_{1}v_{3}\}$ is isomorphic to $G+\{v_{2}v_{3}\}$, and both have broadcast less than $2\alpha^{2\ell}$. Meanwhile, in the graph $G+\{v_{1}v_{3},v_{2}v_{3}\}$, every pair of vertices lies on some cycle of length at most $22\ell+$ 2. Thus, the broadcast of this graph is at least $\alpha^{\ell+1}$. Given a sufficiently large value of $\ell$, we now have
$\beta({G+\{v_{1}v_{3}\}})-\beta({G})<\alpha^{2\ell}<\alpha^{\ell+1}-2\alpha^{2\ell}<\beta({G+\{v_{1}v_{3},v_{2}v_{3}\}})-\beta({G+\{v_{2}v_{3}\}}),$
violating the definition of submodularity. A similar analysis on the same graph can be used to show that the logarithm of broadcast is not submodular. In this case, one need only show that $\frac{\beta({G+\{v_{1}v_{3}\}})}{\beta({G})}<\frac{\beta({G+\{v_{1}v_{3},v_{2}v_{3}\}})}{\beta({G+\{v_{2}v_{3}\}})}$ for an appropriately selected value of $\ell$.
However, we next show that when we restrict the set of candidate edges to edges out of a “center” vertex $u$, submodularity holds for proximities of $u$ to other vertices. Specifically, let $E^{u}$ be $\{uv:v\in V\}$. For a subset $S\subseteq E^{u}$, define
$g_{v}(S):=\log\mathcal{P}_{G+S}({v},{u}).$
$\mathcal{E}_{0}$$\mathcal{E}_{1}$$\mathcal{E}_{2}$$x_{b}$$x_{c}$$x_{d}$
Figure 3: Events and probabilities used in the proof of Lemma 6.
By subtracting a term corresponding to the proximity in $G$ (without any edge additions), $g_{v}$ lets us measure the “gain” in the proximity between $u$ and $v$ provided by adding edges $S$. This will be used crucially in the algorithm. We first prove submodularity and monotonicity:
Lemma 6.
For any graph $G$, vertices $u,v$, the function $g_{v}:2^{E^{u}}\mapsto\mathbb{R}$ is monotone and submodular.
Proof.
Since proximity only increases by adding edges (this fact can readily be seen by the sampling based definition of proximity), it is clear that $g_{v}$ is monotone. So let us focus on submodularity.
We first study the two-edge setting: consider any $e_{1},e_{2}\in E^{u}$. We will claim that
$\frac{\mathcal{P}_{G+e_{1}+e_{2}}({v},{u})}{\mathcal{P}_{G+e_{2}}({v},{u})}\leq\frac{\mathcal{P}_{G+e_{1}}({v},{u})}{\mathcal{P}_{G}({v},{u})}.$ (2)
The inequality is obvious if either $e_{1}$ or $e_{2}$ already exists in $G$, and so let us assume that this is not the case.
Let us define $P_{1}$ (resp., $P_{2}$) to be the set of simple paths in $G+e_{1}$ (resp., $G+e_{2}$) that go from $v$ to $u$, ending in the edge $e_{1}$ (resp., $e_{2}$). Let $P_{0}$ be the set of simple paths in $G$ that go from $v$ to $u$, (so they do not contain either of $e_{1}$ or $e_{2}$). The key observation is that any simple path from $v$ to $u$ in $G+e_{1}+e_{2}$ must be in $P_{0}\cup P_{1}\cup P_{2}$. This is because no simple path contains both $e_{1}$ and $e_{2}$; it also cannot have $e_{1}$ or $e_{2}$ as an intermediate edge in the path. Now, define $\mathcal{E}_{0},\mathcal{E}_{1}$, and $\mathcal{E}_{2}$ to be the events that at least one path from $P_{0}$, $P_{1}$, or $P_{2}$ (respectively) exists in a sampled graph. Thus, the inequality (2) is equivalent to:
$\frac{\Pr[\mathcal{E}_{0}\cup\mathcal{E}_{1}\cup\mathcal{E}_{2}]}{\Pr[\mathcal{E}_{0}\cup\mathcal{E}_{2}]}\leq\frac{\Pr[\mathcal{E}_{0}\cup\mathcal{E}_{1}]}{\Pr[\mathcal{E}_{0}]}.$
To simplify the following algebra, we introduce some variables; see Figure 3. Specifically, we say that $x_{a}=\Pr[\mathcal{E}_{0}]$, $x_{b}=\Pr[\bar{\mathcal{E}_{0}}\cap\mathcal{E}_{1}\cap\bar{\mathcal{E}_{2}}]$, $x_{c}=\Pr[\bar{\mathcal{E}_{0}}\cap\mathcal{E}_{1}\cap\mathcal{E}_{2}]$, and $x_{d}=\Pr[\bar{\mathcal{E}_{0}}\cap\bar{\mathcal{E}_{1}}\cap\mathcal{E}_{2}]$. Using this notation, it is easy to check that
$\frac{\Pr[\mathcal{E}_{0}\cup\mathcal{E}_{1}\cup\mathcal{E}_{2}]}{\Pr[\mathcal{E}_{0}\cup\mathcal{E}_{2}]}=\frac{x_{a}+x_{b}+x_{c}+x_{d}}{x_{a}+x_{c}+x_{d}}\leq\frac{x_{a}+x_{b}+x_{c}}{x_{a}+x_{c}}\leq\frac{x_{a}+x_{b}+x_{c}}{x_{a}}=\frac{\Pr[\mathcal{E}_{0}\cup\mathcal{E}_{1}]}{\Pr[\mathcal{E}_{0}]},$
establishing the claim.
The claim implies submodularity in a straightforward way: suppose $S\subseteq T\subseteq E^{u}$ and consider any $e\in E^{u}$. Again, the case $e\in T$ is trivial, so let us assume that $e\not\in T$. We can now consider $G+S$ to be the “base” graph, and use the argument above repeatedly, choosing $e_{1}=e$, and an element of $T\setminus S$ as $e_{2}$. This gives us
$\frac{\mathcal{P}_{G+T+e}({v},{u})}{\mathcal{P}_{G+T}({v},{u})}\leq\frac{\mathcal{P}_{G+S+e}({v},{u})}{\mathcal{P}_{G+S}({v},{u})}.$
Taking logarithms, we obtain that $g_{v}$ is submodular. ∎
Our algorithm is based on a potential function that captures the proximity between all pairs $(i,j)$ of vertices. Before describing it, we introduce an auxiliary function that is defined for one pair $(i,j)\in V^{2}$, a given vertex $u$, a “target” proximity value $\beta^{\prime}$, and a set of edges $T$ all incident to $u$:
$\mu(i,j;T):=-\log\mathcal{P}_{G+T}({u},{i})-\log\mathcal{P}_{G+T}({u},{j})+\log\frac{\beta^{\prime}}{2}.$
Note that by Lemma 6, for any $(i,j)$ we have that $\mu(i,j;T)$ is submodular. The algorithm assumes as parameters a vertex $u$ (we will need to run the algorithm for every choice of $u$), a parameter $\varepsilon>0$, and a target broadcast value $\beta^{\prime}$. The algorithm is as follows:
1
2
3Initialize $S^{(0)}=\emptyset$, $t=0$
4 Define $A=\{(i,j)\in V^{2}:\mathcal{P}_{G}({i},{j})<\beta^{\prime}/2\}$ (Call these “active” pairs)
5 For any $T$, define $\Psi(T)=\sum_{(i,j)\in A}\max\{0,\mu(i,j;T)\}$
6while $\Psi(S^{(t)})>\log(1/\alpha^{\varepsilon})$ do
7       Find edge $e$ incident to $u$ that minimizes $\Psi(S^{(t-1)}\cup\{e\})$
8       Increment $t$; define $S^{(t)}=S^{(t-1)}\cup\{e\}$
Return $S^{(t)}$
Algorithm 1
While submodularity will ensure that the drop in potential is significant at every step, it turns out that it is non-trivial to prove that the optimal subset achieves low potential! This is indeed the reason for our definition and use of active pairs in the algorithm. The key technical lemma is the following:
Lemma 7.
Let $u$ be the center of the star edges $E^{\prime}$ obtained from Theorem 6 and $\beta^{\prime}$ be the corresponding broadcast value. Suppose $i,j\in V$ such that $\mathcal{P}_{G}({i},{j})<\beta^{\prime}/2$. Then $\mathcal{P}_{G+E^{\prime}}({i},{u})\cdot\mathcal{P}_{G+E^{\prime}}({j},{u})\geq\beta^{\prime}/2$.
Proof.
Consider the set of all paths from $i$ to $j$ in $G+E^{\prime}$. These can be partitioned into $P_{u}$ and $P_{0}$ according to whether or not (respectively) they use the vertex $u$. Thus, $P_{u}$ and $P_{0}$ are disjoint sets with $\Pr[P_{u}\cup P_{0}]\geq\beta^{\prime}$, by assumption. Moreover, because every path in $P_{0}$ also exists in $G$, i.e., these paths do not use any edges from $E^{\prime}$, we have $\Pr[P_{0}]\leq\beta^{\prime}/2$. We now apply Lemma 3, and complete the proof by noting that (by definition of proximity) $\mathcal{P}_{G+E^{\prime}}({i},{u})\cdot\mathcal{P}_{G+E^{\prime}}({j},{u})\geq\Pr[P_{u}[i,u]]\cdot\Pr[P_{u}[u,j]]$. ∎
Lemma 7 immediately implies that for $\Psi$ as defined in Algorithm 1, for the star edges $E^{\prime}$, we have $\Psi(E^{\prime})=0$. This is because for every active pair $(i,j)$, the lemma implies that $\mu(i,j;E^{\prime})\leq 0$. We then have the following guarantee for the algorithm, at every step $t$.
Lemma 8.
Let $S^{(t)}$ be the set of added edges as defined in the algorithm. For any $t\geq 1$, we have $\Psi(S^{(t)})\leq\Psi(S^{(t-1)})\left(1-\frac{1}{2k}\right)$.
Proof.
Let $e_{1},e_{2},\dots,e_{2k}$ be the edges in $E^{\prime}$. Define the current set of active pairs as
$A^{(t-1)}=\{(i,j)\in V^{2}:\mu(i,j;S^{(t-1)})>0\}.$
For any such pair, the first observation is to note that
$\sum_{\ell\in[2k]}\left(\mu(i,j;S^{(t-1)})-\mu(i,j;S^{(t-1)}\cup\{e_{\ell}\})\right)\geq\mu(i,j;S^{(t-1)}).$ (3)
This follows via a standard argument, adding the elements $e_{\ell}$ in some order and using submodularity. For a technical reason, we note that this also implies that
$\sum_{\ell\in[2k]}\left(\mu(i,j;S^{(t-1)})-\max\{0,\mu(i,j;S^{(t-1)}\cup\{e_{\ell}\})\}\right)\geq\mu(i,j;S^{(t-1)}).$ (4)
This follows from (3), because if for some $\ell$, $\mu(i,j;S^{(t-1)}\cup\{e_{\ell}\})<0$, that term in the summation alone is $\geq$ RHS, and we only need to use the fact that every other term is non-negative. If all the $\mu$ are $\geq 0$, then (3) and (4) are identical.
We can sum this over all pairs $(i,j)$ in $A^{(t-1)}$, and noting that the RHS is exactly $\Psi(S^{(t-1)})$, we have by averaging,
$\exists\ell\leavevmode\nobreak\ :\leavevmode\nobreak\ \sum_{(i,j)\in A^{(t-1)}}\left(\mu(i,j;S^{(t-1)})-\max\{0,\mu(i,j;S^{(t-1)}\cup\{e_{\ell}\})\}\right)\geq\frac{1}{2k}\Psi(S^{(t-1)}).$
This implies that $\Psi(S^{(t-1)})-\Psi(S^{(t)})\geq\frac{1}{2k}\Psi(S^{(t-1)})$. The lemma follows by rearranging the terms. ∎
The preceding lemmas give us all the tools we need to analyze Algorithm 1.
Theorem 7.
For any $\varepsilon,\alpha>0$, there exists a polynomial-time algorithm which produces broadcast at least $\frac{\beta^{*}\alpha^{2+\varepsilon}}{(1+\varepsilon)(24k^{2}+6)}$ by adding $O(k\log n)$ edges.
Proof.
We begin by assuming that we have a value $\beta^{*}\geq x\geq\beta^{*}/(1+\varepsilon)$. Having $x$, we set $\beta^{\prime}=\frac{x\alpha^{2}}{12k^{2}+3}$, so Theorem 6 tells us that it is possible to achieve broadcast $\beta^{\prime}$ by adding $O(k)$ edges incident to a single vertex. We now proceed with Algorithm 1, trying each possible vertex $u$. Because we may assume connectivity of $G$ (see Section 2), we have $\mathcal{P}_{G}({u},{i})\geq\alpha^{n}$ for every $u,i$. Thus, $\mu(i,j;S^{(0)})\leq\frac{2n}{\varepsilon}\log(1/\alpha^{\varepsilon})$ for all $i,j$, implying that the initial potential is at most $2n^{3}/\varepsilon$. Since the potential drops by a factor at least $(1-\frac{1}{2k})$ in each iteration, and since the algorithm terminates when the potential reaches $\log\frac{1}{\alpha^{\varepsilon}}$, we conclude that the number of iterations is $O(k\log n)$. Furthermore, when the algorithm terminates, we have $\mu(i,j;S^{(t)})\leq\log\frac{1}{\alpha^{\varepsilon}}$ for all pairs $i,j$, this implies that for all $(i,j)$, $\mathcal{P}_{G+S^{(t)}}({i},{j})\geq\frac{\beta^{\prime}}{2}\cdot\alpha^{\varepsilon}\geq\frac{\beta^{*}\alpha^{2+\varepsilon}}{(1+\varepsilon)(24k^{2}+6)}$.
To obtain the estimate $x$ for $\beta^{*}$, we use a technique similar to that of Theorem 5. Specifically, we note that a standard analysis can be used to obtain a precise bound $b\in O(k\log n)$ on the number of edges added by our algorithm, given that $\beta^{*}\geq x$ and that we have guessed the correct vertex $u$. We conduct a binary search of integers in the interval $[0,\log_{1+\varepsilon}\frac{1}{\beta({G})}]$. For each such integer $i$, we execute our algorithm with $x=\beta({G})(1+\varepsilon)^{i}$. If the algorithm adds more than $b$ edges for every guess of the vertex $u$, then we conclude that $x>\beta^{*}$. Let $i^{*}$ be the largest guessed value such that the algorithm terminates (for some guess of vertex $u$) after adding at most $b$ edges. Then, as previously argued in the proof of Theorem 5, we have that $x=\beta({G})(1+\varepsilon)^{i^{*}}\geq\beta^{*}/(1+\varepsilon)$. This completes the proof. ∎
6 Hardness of Broadcast Improvement
In this section we establish hardness of the Broadcast Improvement problem by presenting lower bounds for bicriteria approximation algorithms.
Theorem 8.
For any constants $c^{\prime}\geq 1$ and $\varepsilon>0$, unless P=NP, there is no polynomial-time algorithm which can guarantee broadcast at least $(\beta^{*}{})^{\frac{6}{5}-\varepsilon}$ while adding at most $c^{\prime}k$ edges.
The reduction is from a variant of the Set Cover problem. Specifically, we rely on the following hardness assumption  [12, 14, 13]:
Assumption 1.
[Gap Set Cover] Let $c\geq 1$ be any constant. Given a collection of $n$ sets $S_{1},S_{2},\dots,S_{n}\subseteq[m]$, it is NP-hard to distinguish between two following cases:
  • Yes: There are $k$ sets in the collection whose union is $[m]$.
  • No: The union of any $ck$ sets can cover at most $(1-(1-\frac{1}{k})^{ck}+\delta)m$ elements (where $\delta>0$ is any small enough constant that keeps the term in the parentheses $<1$).
Furthermore, the hardness holds even when $m=\Theta(n)$ and $|S_{i}|=O(\text{polylog}(n))$.
Note that the assumption implies that doing even slightly better than the bicriteria guarantee of the greedy algorithm of Section 5 is NP-hard. For our reduction, we only need a bound of $(1-\Omega(1))m$ in the No case, which is weaker. Likewise, we only require $|S_{i}|=o(\sqrt{n})$. We also remark that Assumption 1 likely holds even with $c=(\log n)^{1-o(1)}$. With this stronger assumption, our hardness results can be improved to nearly match our algorithmic bounds. We omit the details.
Proof.
Our reduction from Gap Set Cover is as follows.
Instance: Given an instance of Gap Set Cover consisting of a collection of $n$ sets $S_{1},S_{2},\dots,S_{n}\subseteq[m]$. We construct an Broadcast Improvement instance of $(G=(V,E,\alpha),k)$ as follows:
We create a graph $G$ with a pivot vertex $p$, vertices $s_{i}$ corresponding to sets $S_{i}$ (called set vertices) and vertices $e_{i}$ corresponding to elements $i\in[m]$ (called element vertices). Between every pair of set vertices $s_{i},s_{j}$, we add a path of length $l$, where $l$ is an even integer parameter whose value will be specified later. These paths are mutually disjoint, and so there are $\binom{n}{2}(l-1)$ vertices along the paths. We call these set-set internal vertices. Next, we add a path of length $l$ between $s_{i}$ and $e_{j}$ for all $j\in S_{i}$. (I.e., we connect a set vertex $s_{i}$ to all the element vertices $e_{j}$ corresponding to elements $j\in S_{i}$.). Once again, these paths are all mutually disjoint. We call the vertices on the paths the set-element internal vertices. Finally, we connect the pivot to each set vertex via mutually disjoint paths of length $l$. We call the internal vertices along these paths pivot-set internal vertices.
Now we argue about the maximum broadcast that can be achieved after adding $k$ edges to such a graph $G$.
Figure 4: An example instance of Broadcast Improvement constructed by our hardness reduction, after adding $k$ edges. We use $s_{i}$ for set vertices, $e_{i}$ for element vertices, and dot-dash lines to indicate the $l$-length paths joining these. Internal vertices which are the endpoint of a new edge are shown as smaller black vertices along these paths. Among the $k$ added edges (solid), we highlight one of most types of possible edge additions as $f_{1},\ldots,f_{5}$, and indicate using the associated color the set and element vertices they involve. For clarity, several types of edge additions incident to $p$ are omitted, along with the $l$-length paths between $p$ and each set vertex. Finally, $e_{i}$ and $e_{j}$ are a pair of element vertices outside the involved set that have a distance of at least $3l$ between them. These are the vertices that we use to prove (in the no-case) that the broadcast is at most $\alpha^{3l-\varepsilon}$.
Yes-case: let there be a set cover of size $k$ that covers all elements. Suppose we consider adding $k$ edges between the pivot vertex $p$ and set vertices corresponding to sets in the set cover. We claim that between any two vertices in the resulting graph, there is a path of length at most $12.5l+$ 1, thus implying that the broadcast $\beta({G})\geq\alpha^{2.5l+1}$.
To see the claim, we argue separately for each vertex:
First, from the pivot vertex $p$, every vertex can be reached via a path of length $\leq 2l$. To see this, note that the distance of each $e_{i}$ (and therefore also each set-element internal vertex) from $p$ is at most $l+1$; this follows because we have direct edges from $p$ to a set cover. Also, every set-vertex (and therefore also every pivot-set internal vertex) can be reached via a path of length at most $l$. Thus, all of the set-set internal vertices can be reached from $p$ via a path of length $2l$ (indeed, this can be made $1.5l$ by choosing the closer set-vertex).
Second, from any set vertex, we can reach every other set vertex using a path of length $l$, and thus every element vertex with a path of length $\leq 2l$. Further, any of the internal vertices can also be reached via a path of length $\leq 2l$, as can the pivot.
Third, from any element vertex $e_{i}$, we can reach every other element vertex $e_{j}$ with a path of length $\leq 2l+2$ (by going to the set vertex covering $i$, going to the pivot, then to the set vertex covering $j$, then going to $e_{j}$). Further, any set vertex can be reached via a path of length at most $2l$ (going to a set vertex covering $i$ and taking the length $l$ path to the desired set vertex). Any set-set internal vertex can thus be reached via a path of length at most $2.5l$: we can go from the target vertex to the closest set vertex —with a path of length $\leq 0.5l$— and from there to $e_{i}$ by a path of length $\leq 2l$ as before. Moreover, any pivot-set internal vertex can be reached via a path of length at most $2l$, using a path of length at most $l+1$ to the pivot and then proceeding to the target vertex via path of length at most $l-1$. Finally, any set-element internal vertex can be reached via a path of length $(2.5l+1)$: from the target, we can either go to an element vertex via a path of length $(0.5l-1)$ or a set vertex via a path of length $(0.5l+1)$, and using the above, this implies that we can get to the target by a path of length $12.5l+$ 1.
Fourth, from any pivot-set internal vertex $y$, we can reach any other pivot-set internal vertex via a path of length $<2l$ by first traveling to the pivot, and then to the target vertex. Moreover, we can reach any set-element internal vertex $x$ via a path of length at most $2.5l$. We begin by traveling to either the pivot or the set vertex corresponding to $y$ (whichever is closer) via a path of length at most $0.5l$. We then continue to $x$ via at most $2l$ additional edges, as argued above.
Next, from any set-set internal vertex, we can reach any set vertex with a path of length $\leq 1.5l$, and thus we can reach every other vertex with a path of length $\leq 2.5l$.
Finally, from a set-element internal vertex, it only remains to show that we can reach any other set-element internal vertex using a short path (other cases are covered above by symmetry). Consider two set-element internal vertices $x$ and $x^{\prime}$, with $x$ (resp., $x^{\prime}$) on the path from $s_{i}$ to $e_{j}$ (resp., $s_{i^{\prime}}$ to $e_{j^{\prime}}$). We show that there is a cycle with $25l+$ 2 edges in the graph that contains $x,x^{\prime}$. Note that this implies that there is a path of length $\leq 2.5l+1$ (i.e., the shorter path on the cycle). As shown in Figure 5, note that there is a path of length $22l+$ 2 between $e_{j}$ and $e_{j^{\prime}}$ (using the covering set vertices, as seen above), and there is also the path going to $s_{i}$, to $s_{i^{\prime}}$, then to $e_{j^{\prime}}$.
Figure 5: This figure shows the cycle of length $25l+$ 2 containing two set-element internal vertices $x$ and $x^{\prime}$. This cycle shows that $x$ and $x^{\prime}$ have proximity at least $\alpha^{2.5l+1}$ (in the yes-case).
This completes the proof of the claim. Thus, in the yes-case, $\beta({G})\geq\alpha^{2.5l+1}$.
No-case: By 1, the union of any $ck$ sets among $S_{1},\dots,S_{n}$ covers at most $(1-\Omega(1))m$ elements, for any constant $c\geq 1$.
Now consider adding $r=c^{\prime}k$ edges, where $r\leq ck/8$. Let $E^{\prime}$ denote the set of added edges and let $G^{\prime}$ be the graph obtained from $G$ by adding the edges in $E^{\prime}$.
For any $\{u,v\}\in E^{\prime}$, we define a set of “involved” set and element vertices as follows. A set vertex $s_{i}$ is said to be involved in edge $\{u,v\}$ if for $x\in\{u,v\}$, we have (a) $s_{i}=x$, (b) $x$ is a set-set internal vertex and one of the end-points of the corresponding path (the one containing $x$) is $s_{i}$, or (c) $x$ is a set-element internal vertex and the set end-point of the corresponding path is $s_{i}$. Analogously, we say that an element vertex $e_{i}$ is involved in edge $\{u,v\}$ if for $x\in\{u,v\}$, either (a) $e_{i}=x$ or (b) $x$ is a set-element internal vertex and $e_{i}$ is the element end-point of the corresponding path.
We also generalize the notation slightly and say that a vertex is involved in a set of edges $E^{\prime}$ if it is involved in at least one of the edges $e\in E^{\prime}$. The main claim due to our choice of parameters is the following.
Claim.
There exist element vertices $e_{i},e_{j}$ such that (a) neither of them is involved in $E^{\prime}$, (b) none of the sets containing $i,j$ are involved in $E^{\prime}$, and (c) there is no set $S_{t}$ that contains both $i$ and $j$.
Proof of Claim..
First, note that any edge $e\in E^{\prime}$ can have at most $4$ set vertices and $2$ element vertices involved in it. Thus, if we pick $r$ edges, we will have at most $4r$ set vertices and $2r$ element vertices involved. For each of the involved element vertices, choose an arbitrary set that covers that element, thus obtaining a set $I$ of $6r$ set vertices with the property that all the set and element vertices involved in $S^{\prime}$ are either in $I$ or are covered by $I$. Now by our assumption for the no-case and the choice of $r$, this means there are $\Omega(m)$ element vertices that are not covered by the sets in $I$; Call this set $U$. By definition, for any $e_{i}\in U$, none of the sets covering $e_{i}$ are in $I$ (and so property (b) automatically holds). Finally, note that there must exist $e_{i},e_{j}\in U$ such that no set vertex has length $l$ paths to both of them. This can be seen by a simple averaging argument as follows. Let $c_{ij}$ be the number of sets that contain both $i$ and $j$. Then, we have
$\sum_{i,j\in U}c_{ij}\leq\sum_{\ell\in[n]}\binom{|S_{\ell}|}{2}\leq n\cdot\text{polylog}(n).$
Indeed, if we take all pairs $i,j\in[m]$, the first inequality becomes equality. Now, if $c_{ij}\geq 1$ for all $i,j\in U$, then $\sum\limits_{i,j\in U}c_{ij}\geq\binom{|U|}{2}$, and since $|U|=\Omega(m)$ and $m=\Theta(n)$, this leads to a contradiction. Thus, there must exist $i,j\in U$ such that $c_{ij}=0$, and this completes the proof of part (c) of the claim. ∎
For $e_{i},e_{j}$ satisfying the claim, we show that $\mathcal{P}_{G^{\prime}}({e_{i}},{e_{j}})$ has to be small. First, note that the distance between them in $G^{\prime}$ is $\geq 3l$. This is because any path from $e_{i}$ to $e_{j}$ must go via one of the set vertices containing $i,j$, and since they are not involved in $E^{\prime}$, the shortest path between those set vertices has length $l$. Our goal now is to argue that the broadcast is also “close” to $\alpha^{3l}$. For this, we show how to simplify the graph for easier reasoning about broadcast.
Claim.
[Subset contraction] Let $S$ be any subset of the vertices of $G^{\prime}$. Define a contraction as the process where we replace $S$ with a single “hub” vertex $h$, and replace every edge of the form $\{u,v\}$ where $u\in S$ and $v\not\in S$ with $\{h,v\}$ (forming parallel edges if appropriate). Let $G^{c}$ be the graph obtained after contraction. For any $u,v\not\in S$, we have $\mathcal{P}_{G^{c}}({u},{v})\geq\mathcal{P}_{G^{\prime}}({u},{v})$.
The claim then follows immediately from the sampling-based definition of proximity: suppose we sample edges with probability $\alpha$ each, then if a path exists in $G^{\prime}$, it also exists in $G^{c}$ (because of us placing parallel paths). Now given $E^{\prime}$, define $S^{\prime}$ as the set of all set and element vertices involved in $E^{\prime}$, along with the pivot $p$. Then, define $S$ to be the union of $S^{\prime}$ and all the internal vertices along paths between vertices of $S^{\prime}$. The crucial observation now is that every edge in $E^{\prime}$ has both its end-points in $S$.
Now suppose we contract the set $S$ in $G^{\prime}$ and obtain the graph $G^{c}$. By the claim, it suffices to show an upper bound on $\mathcal{P}_{G^{c}}({e_{i}},{e_{j}})$ (where $e_{i},e_{j}$ are the element vertices that we identified earlier). To do this, we make another observation about $G^{c}$: its vertices consist of $h$ (the new hub vertex), a subset $V_{set}$ of the original set vertices, a subset $V_{elt}$ of the element vertices, all the internal vertices of the paths between $V_{set}\cup V_{elt}$, and all the internal vertices of the paths between $h$ and $V_{set}\cup V_{elt}$. Thus it is natural to define a “path compressed” graph $H$, whose vertex set is $\{h\}\cup V_{set}\cup V_{elt}$, which has an edge iff there is a path of length $l$ in $G^{c}$. Note that there can be parallel edges in $H$. Now suppose we view $H$ as an information graph, where the sampling probability is $\alpha^{l}$ for every edge. Then we have the following easy observation:
Observation 2.
For all $u,v\in\{h\}\cup V_{set}\cup V_{elt}$, we have $\mathcal{P}_{G^{c}}({u},{v})=\mathcal{P}_{H}({u},{v})$.
For our $e_{i},e_{j}$ of interest, bounding $\mathcal{P}_{H}({e_{i}},{e_{j}})$ will turn out to be simple, because the edge probability $\alpha^{l}$ will be chosen to be so small that only the shortest path between $e_{i}$ and $e_{j}$ matters. We formalize this in the following simple claim.
Claim.
Let $H=(V_{H},E_{H})$ be an information graph on $n_{H}$ vertices in which every edge has sampling probability $\leq\delta<\frac{1}{2n_{H}}$. Let $u,v\in V$ such that dist$(u,v)=c$, for some integer $c\geq 1$. Then $\mathcal{P}_{H}({u},{v})\leq 2n_{H}^{c-1}\delta^{c}$.
Proof of Claim..
For any path of length $\ell$, the probability that the path exists in a sampled graph is $\delta^{\ell}$. Between any two vertices, there are clearly at most $n_{H}^{\ell-1}$ paths of length $\ell$, and thus by a union bound, noting that dist$(u,v)=c$ and therefore there are no paths of length $<c$, we have:
$\mathcal{P}_{H}({u},{v})$ $\leq\sum_{\ell=c}^{n_{H}-1}n_{H}^{\ell-1}\delta^{\ell}\leq n_{H}^{c-1}\delta^{c}\cdot\sum_{\ell\geq 0}(\delta n_{H})^{\ell}\leq 2n_{H}^{c-1}\delta^{c}.$
The claim now implies a bound on $\mathcal{P}_{H}({e_{i}},{e_{j}})$; we cannot use it directly since our $H$ has parallel edges, but we note that there are at most $16k+$ 1 parallel edges between any two vertices (because that is a bound on the number of “non-internal” vertices, plus the pivot, in the contracted set $S$). Thus, we can replace parallel edges by a single edge with $\delta\leq(6k+1)\alpha^{l}$, and use 2 to obtain:
$\mathcal{P}_{H}({e_{i}},{e_{j}})=\mathcal{P}_{G^{\prime}}({e_{i}},{e_{j}})\leq\mathcal{P}_{G^{c}}({e_{i}},{e_{j}})\leq 2(m+n)^{2}\delta^{3}.$
This is because the shortest path in $H$ between $e_{i}$ and $e_{j}$ has length three, as there is no set that contains both $i$ and $j$. Now, if we choose $l$ large enough (approximately $\frac{\log n}{\epsilon}$), we can make $2(m+n)^{2}\delta^{3}$ to be $<\alpha^{3l-\epsilon}$, for any $\epsilon>0$.
Since the broadcast in the yes-case is $\geq\alpha^{2.5l+1}$, the desired gap follows.
7 Single-Source Variant
In this section we study the problem of adding edges to improve the information access of a single vertex $v_{s}$. That is, we seek to maximize the reach $\Upsilon_{G}({v_{s}})$ of $v_{s}$.
Reach Improvement
Input: An information graph $G=(V,E,\alpha)$, a vertex $v_{s}\in V$, and a non-negative integer $k$. Task: Add $k$ edges to $G$ such that the reach of $v_{s}$ in the resulting information graph is maximized.
We first show that the Broadcast Improvement algorithms which we developed in Sections 3-5 also apply in this context, with constrained loss in the guarantees.
Lemma 9.
Suppose for some functions $f,g$ a polynomial-time algorithm exists which produces broadcast at least $f(\beta^{*}{},k,\alpha)$ using $g(k)$ edge additions, given any instance of Broadcast Improvement. Then a polynomial-time algorithm exists which produces reach at least $f((\Upsilon^{*}{})^{2},k,\alpha)$ using $g(k)$ edge additions, given any instance of Reach Improvement.
Proof.
Let $G=(V,E,\alpha)$ be an information graph and $v_{s}\in V$ a source vertex. Furthermore, suppose that $\beta^{*}$ is the maximum broadcast achievable by adding $k$ edges to $G$, and that $\Upsilon^{*}$ is the maximum achievable reach of $v_{s}$ with $k$ edge additions. Suppose toward a contradiction that $\Upsilon^{*}>\sqrt{\beta^{*}}$, and let $G^{*}$ denote the information graph resulting from some optimal set of edge additions for the Reach Improvement problem. Then we have that
$\mathcal{P}_{G^{*}}({u},{v})\geq\mathcal{P}_{G^{*}}({v_{s}},{u})\cdot\mathcal{P}_{G^{*}}({v_{s}},{v})\geq\Upsilon^{*}\cdot\Upsilon^{*}>\beta^{*}$
Because $u$ and $v$ were chosen arbitrarily, we conclude that $\beta({G^{*}})>\beta^{*}$, a contradiction. Now that we have proven $\Upsilon^{*}\leq\sqrt{\beta^{*}}$, it is a straightforward observation that any $g(k)$ edge additions which achieve broadcast at least $f(\beta^{*},k,\alpha)$ also produce reach at least $f(\Upsilon^{*}{}^{2},k,\alpha)$ for every choice of source vertex. ∎
Paying an extra factor two in the exponent is far from desirable. Fortunately, we can use the ideas developed in Section 4 to directly obtain a linear approximation for Reach Improvement using $O(k\log n)$ edges. In the following, we adapt Definition 2 to this setting in the natural way. That is, a set $W_{u}$ is a $(u,r)$-witness if $\mathcal{P}_{G+W_{u}}({v_{s}},{u})\geq r$ and a solution $S$ is $(c,r)$-witnessing if it contains as a subset a $(u,r)$-witness of size at most $c$ for every vertex $u$.
Lemma 10.
For any instance of Reach Improvement, there exists a $(1,\frac{\Upsilon^{*}{}}{2k+2})$-witnessing solution of size at most $2k$.
Proof.
Let $(G=(V,E,\alpha),v_{s},k)$ be an instance of Reach Improvement, and suppose that $S^{*}$ is an optimal solution, so $\Upsilon_{G+S^{*}}({v_{s}})=\Upsilon^{*}$. We construct an alternate solution $S$ as follows. First, impose an arbitrary ordering on the endpoints of the edges in $S$: $s_{1},s_{2},\ldots s_{m}$, and we set $S=\{v_{s}s_{1},v_{s}s_{2},\ldots v_{s}s_{m}\}$. It is clear that $S$ has size at most $2k$, so what remains is to prove that $S$ contains a $(u,\frac{\Upsilon^{*}}{2k+2})$-witness of size at most 1 for every $u\in V$.
Let $P_{v_{s}u}$ be all paths from $v_{s}$ to $u$ in $G+S^{*}$. We partition these paths according to the last edge in $S^{*}$ to appear along the path. This creates at most $k+1$ equivalence classes, including the class of paths which already existed in $G$ before any edge additions were made. The sum of the contributions of each equivalence class is an upper bound for $\mathcal{P}_{G+S^{*}}({v_{s}},{u})$, and therefore also for $\Upsilon^{*}$. Then there is at least one equivalence class with contribution at least $\Upsilon^{*}/(k+1)$. If one such class is the set of paths using no edges from $S^{*}$, then the empty set is our witness. Otherwise, let $P_{v_{s}u}^{ij}$ be such an equivalence class, corresponding to edge $s_{i}s_{j}\in S^{*}$. We further partition this set of paths according to which of $s_{i}$ or $s_{j}$ is the trailing vertex of $s_{i}s_{j}$ in a given path. One of these subsets has contribution at least half of that of $P_{v_{s}u}^{ij}$, i.e., at least $\Upsilon^{*}/(2k+2)$. Call this subset $\mathcal{C}$, and assume without loss of generality that it is the subset corresponding to vertex $s_{i}$.
We now show how to replace $\mathcal{C}$ with a new set of paths $Q$ which has contribution at least $\Upsilon^{*}/(2k+2)$ and which contains only one edge, specifically $v_{s}s_{i}$, from $S$. This completes the proof, as it shows that $\{v_{s}s_{i}\}$ is a $(u,\frac{\Upsilon^{*}}{2k+2})$-witness of size 1. For each path $p\in\mathcal{C}$, we replace $p[v_{s},s_{i}]$ with the edge $v_{s}s_{i}$. By construction, $v_{s}s_{i}$ is the only edge from $S$ contained in any path in $Q$. Moreover, we note that $Q[s_{i},u]=\mathcal{C}[s_{i},u]$. Let $\mathcal{E}_{1}$ be the event that a path from one of these segments is sampled. We note that for a path in $\mathcal{C}$ to be sampled, we require both event $\mathcal{E}_{1}$ and that the edge $s_{j}s_{i}$, which appears on no path in $\mathcal{C}[s_{i},u]$, is sampled. Thus, we have $\Pr[\mathcal{C}]\leq\alpha\Pr[\mathcal{E}_{1}]$. Meanwhile, $\Pr[Q]=\alpha\Pr[\mathcal{E}_{1}]$, as the paths in $Q[s_{1},u]$ and the edge $v_{s}s_{i}$ are sampled independently. The result follows.
We now show how to use $O(k\log n)$ edges to obtain a guarantee of either $\frac{\Upsilon^{*}}{(1+\varepsilon)(2k+2)}$ or $\frac{\Upsilon^{*}\alpha}{(1+\varepsilon)(1+2k\alpha)}$. Both algorithms work via reductions to Hitting Set, with the former guarantee relying on Lemma 10 and the latter on Lemma 1. The former guarantee is stronger if $\alpha<\frac{1}{2}$ and the latter is stronger if $\alpha>\frac{1}{2}$. Similarly to Theorem 5, we also use the reductions to Hitting Set to conduct a binary search until we have found a good estimate for the value of $\Upsilon^{*}$.
Theorem 9.
Given any $\alpha,\varepsilon>0$, there exists a polynomial-time algorithm which produces reach (for a given source vertex $v_{s}$) at least $\max\{\frac{\Upsilon^{*}}{(1+\varepsilon)(2k+2)},\frac{\Upsilon^{*}\alpha}{(1+\varepsilon)(1+2k\alpha)}\}$ using $O(k\log n)$ edge additions.
Proof.
Let $(G=(V,E,\alpha),v_{s},k)$ be an instance of Reach Improvement. We begin by assuming that we already know the optimum achievable reach $\Upsilon^{*}$. We reduce to Hitting Set. If $\alpha\leq\frac{1}{2}$, then the reduction proceeds as follows. The elements of our Hitting Set instance are $V^{2}\setminus E$, i.e., all possible edge additions. For each vertex $u\in V$ with $\mathcal{P}_{G}({v_{s}},{u})<\Upsilon^{*}$, we create a set $\mathcal{W}_{u}$ consisting of all single edge-additions which improve the proximity of $v_{s}$ to $u$ to at least $\frac{\Upsilon^{*}}{2k+2}$. That is, $\mathcal{W}_{u}$ is the set of all $(u,\frac{\Upsilon^{*}}{2k+2})$-witnesses of size 1. According to Lemma 10, there exists a hitting set of size at most $2k$. We use the well-known greedy approximation for Hitting Set [24] to obtain a hitting set of size $O(k\log n)$, and we return these edges as our solution. It follows from the construction that this solution achieves reach at least $\frac{\Upsilon^{*}}{2k+2}$.
On the other hand, if $\alpha>\frac{1}{2}$, then we reduce to Hitting Set as follows. Our elements are $V\setminus v_{s}$, i.e., all vertices except the source. For each vertex $u\neq v_{s}$, we add a set equal to the $\frac{\Upsilon^{*}}{1+2k\alpha}$-neighborhood of $u$. We claim that there exists a hitting set of size at most $2k$. Otherwise, for any optimal solution $S^{*}$ which adds $k$ edges, there exists some vertex $u\neq v_{s}$ such that the endpoints of these edges are disjoint from the $\frac{\Upsilon^{*}}{1+2k\alpha}$-neighborhood of $u$. But then Lemma 1 gives us that $\mathcal{P}_{G+S^{*}}({u},{v_{s}})<\Upsilon^{*}$, a contradiction. Hence, we can once again use the greedy approximation for Hitting Set to obtain a set of $O(k\log n)$ vertices. We add the edges between these vertices and $v_{s}$, guaranteeing reach at least $\frac{\Upsilon^{*}\alpha}{1+2k\alpha}$ as desired.
It remains to show how we can estimate $\Upsilon^{*}$. In the following, let $b$ be the precise bound on edge additions given by the algorithm described above. That is, $b$ is $2k$ multiplied by the precise factor given by the algorithm in [24]. We note that $\Upsilon_{G}({v_{s}})\leq\Upsilon^{*}\leq 1$, so there exists an integer $0\leq i\leq\log_{i+\varepsilon}\frac{1}{\Upsilon_{G}({v_{s}})}$ such that $\Upsilon_{G}({v_{s}})(1+\varepsilon)^{i}\leq\Upsilon^{*}\leq\Upsilon_{G}({v_{s}})(1+\varepsilon)^{i+1}$. We conduct a binary search of integers in the interval $[0,\log_{i+\varepsilon}\frac{1}{\Upsilon_{G}({v_{s}})}]$. For each tested value $j$, we assume that $\Upsilon^{*}=\Upsilon_{G}({v_{s}})(1+\varepsilon)^{j}$ and run the algorithm described above. If the resulting set of edge additions has size greater than $b$, then we use Lemma 10 to conclude that $j>i$. Now, let $j^{*}$ be the largest integer in the interval for which the algorithm adds at most $b$ edges. Then we can conclude that $\Upsilon^{*}\leq\Upsilon_{G}({v_{s}})(1+\varepsilon)^{j^{*}+1}$, and the algorithm gives us a set of at most $b\in O(k\log n)$ edge additions which achieve reach at least
$\frac{\Upsilon_{G}({v_{s}})(1+\varepsilon)^{j^{*}}}{2k+2}=\frac{\Upsilon_{G}({v_{s}})(1+\varepsilon)^{j^{*}+1}}{(1+\varepsilon)(2k+2)}\geq\frac{\Upsilon^{*}}{(1+\varepsilon)(2k+2)}.$
So we have successfully computed the desired approximation. ∎
We conclude this section by showing that no polynomial time algorithm using $O(k)$ edge additions can achieve a linear approximation for Reach Improvement, unless P = NP.
Theorem 10.
For any $\varepsilon,\alpha>0$ and any constant $c\geq 1$, it is NP-hard to provide reach at least $\big{(}\Upsilon^{*})^{\frac{4}{3}-\varepsilon}$ using $ck$ edge additions.
Proof.
We once again reduce from Gap Set Cover. Let $(S_{1},S_{2},\ldots S_{n}\subseteq[m],k)$ be an instance of this problem. Recall that it is NP-hard to distinguish between the cases (i) there exists a collection of at most $k$ sets which cover every element, and (ii) any collection of $2ck$ sets leaves at least $12ck+$ 1 elements uncovered888 We know from [12, 14, 13] that it is NP-hard to distinguish between the existence of a set cover with size $k$ and the non-existence of any set cover with size $2ck$. Our stronger hardness assumption can be obtained via a simple reduction: copy the instance of Gap Set Cover and create $2ck$ additional replicas of each element, giving each replica the same set memberships as the original.  [12, 14, 13]. Let $d$ be the maximum size of any set, i.e., $d=\max_{i\in[n]}|S_{i}|$. Let $b$ be the maximum number of sets containing any single element, i.e., $b=\max_{j\in[m]}|\{i\in[n]\colon j\in S_{i}\}|$. Finally, we will need an additional value $l$, which can be thought of as an integer which is polynomial in $m+n$. We will show how to select the exact value of $l$ at the end of the proof.
We construct an instance of Reach Improvement as follows. First we introduce a vertex $v_{s}$, which will be the source vertex of our constructed instance. Next, for each set $S_{i}$, $i\in[n]$, we introduce a vertex $v_{i}$. We call these vertices set vertices. For each set vertex $v_{i}$, we introduce $l-1$ auxiliary vertices and $l$ edges, such that $v_{i}$ is connected to $v_{s}$ by a path of edge-length $l$ on these vertices and edges. We call these auxiliary vertices the set-path vertices corresponding to set $S_{i}$. So far, we have added (less than) $1+ln$ vertices and exactly $ln$ edges. Next, for each element $j\in[m]$, we add a new vertex $v_{j}$, which we call an element vertex. For each set $S_{i}$ containing element $j$, we add $l-1$ auxiliary vertices and $l$ edges such that $v_{i}$ and $v_{j}$ are connected by a path of edge-length $l$ on these vertices and edges. We call these auxiliary vertices the set-element-path vertices corresponding to set $S_{i}$ and element $j$. This step adds at most $m\cdot d\cdot l$ vertices and edges. We call the information graph we have constructed $G$. We set the limit on edge additions to $k$, and the target value for the reach of $v_{s}$ to $\alpha^{\frac{3l}{2}+1}$. This completes the construction of our instance of Reach Improvement.
It remains to show that the reduction is correct. We begin by assuming that ($(\{S_{i}\}_{i\in[n]},k)$ is a yes-instance. That is, we assume that there exist $k$ sets $S_{i_{1}},S_{i_{2}},\ldots,S_{i_{k}}$ which cover every element. In this case, we add the $k$ edges between $v_{s}$ and the set vertices corresponding to this set cover. That is, we propose the solution $S=\{v_{s}v_{i_{1}},v_{s}v_{i_{2}},\ldots,v_{s}v_{i_{k}}\}$. Now we show that $\Upsilon_{G+S}({v_{s}})\geq\alpha^{\frac{3l}{2}+1}$. Let $u\neq v_{s}$ be some vertex in $G$. If $u$ is a set vertex or a set-path vertex, then our initial construction guarantees that $\mathcal{P}_{G+S}({u},{v_{s}})\geq\mathcal{P}_{G}({u},{v_{s}})\geq\alpha^{l}$. If $u$ is an element vertex, then we identify a set $S_{i}$ which contains the element corresponding to $u$ and is part of the cover. Then
$\mathcal{P}_{G+S}({u},{v_{s}})\geq\mathcal{P}_{G+S}({u},{S_{i}})\cdot\mathcal{P}_{G+S}({S_{i}},{v_{s}})\geq\alpha^{l}\cdot\alpha=\alpha^{l+1}$
Finally, assume that $u$ is a set-element-path vertex corresponding to set $S_{i}$ and element $j$. Observe that either $\mathcal{P}_{G}({u},{v_{i}})\geq\alpha^{\frac{l}{2}}$ or $\mathcal{P}_{G}({u},{v_{j}})\geq\alpha^{\frac{l}{2}}$. In the former case,
$\mathcal{P}_{G+S}({u},{v_{s}})\geq\alpha^{\frac{l}{2}}\cdot\mathcal{P}_{G+S}({v_{i}},{v_{s}})\geq\alpha^{\frac{3l}{2}}$
and in the latter case
$\mathcal{P}_{G+S}({u},{v_{s}})\geq\alpha^{\frac{l}{2}}\cdot\mathcal{P}_{G+S}({v_{j}},{v_{s}})\geq\alpha^{\frac{3l}{2}+1}$
Thus, all vertices have proximity at least $\alpha^{\frac{3l}{2}+1}$ to $v_{s}$ in $G+S$, so $S$ witnesses that $(G,v_{s},k,\alpha^{\frac{3l}{2}+1})$ is a yes-instance of Reach Improvement.
We now assume that ($(\{S_{i}\}_{i\in[n]},k)$ is a no-instance of Gap Set Cover. In this case, we let $S^{*}$ be an optimal solution to our constructed instance of Reach Improvement, and additionally allow that $S^{*}$ may contain up to $ck$ edges. That is, $S^{*}$ is a set of at most $ck$ edge additions, with $\Upsilon_{G+S^{*}}({v_{s}})\geq\Upsilon^{*}$. We will first give an upper bound on $\Upsilon^{*}$, and then show how we could have chosen $l$ such that this upper bound yields the desired hardness result. We impose an arbitrary order on the (at most $2ck$) endpoints of the edges in $S^{*}$, $s_{1},s_{2},\ldots s_{ck}$. We then introduce a new solution $S$ of size at most $2ck$ such that $S=\{v_{s}s_{1},v_{s}s_{2},\ldots,v_{s}s_{2ck}\}$. By carefully inspecting the proof of Lemma 10, we observe that $\Upsilon_{G+S}({v_{s}})\geq\frac{\Upsilon^{*}}{2ck+2}$. We call the vertices $s_{1},s_{2},\ldots s_{2ck}$ the destinations of the solution $S$, and we say that a set $S_{i}$ is involved in solution $S$ if the destinations of $S$ include the set vertex $v_{i}$, any set-path vertex corresponding to $S_{i}$, or any set-element-path vertex corresponding to $S_{i}$. Note that every set-path vertex and every set-element-path vertex corresponds to exactly one set, so since $S$ has at most $2ck$ destinations we can conclude that at most $2ck$ sets are involved in $S$. Next, we say that an element $j$ is uncovered if it is not contained in any set which is involved in $S$. Because only $2ck$ sets are involved in $S$, at least $12ck+$ 1 elements are uncovered. Moreover, because $S$ has at most $2ck$ destinations, there is at least one uncovered element $j$ for which the corresponding element vertex $v_{j}$ is not itself a destination of $S$. We will now show that $\mathcal{P}_{G+S}({v_{j}},{v_{s}})\leq\alpha^{2l}(b+bd)$.
Every path from $v_{j}$ to $v_{s}$ begins with $l$ edges from $v_{j}$ to some set vertex $v_{i}$, where $j\in S_{i}$ and $S_{i}$ is not involved in $S$. From there, paths extend either via $l$ more edges to $v_{s}$, or via $l$ more edges to another element vertex. Each path of the former variety has contribution $\alpha^{2l}$, and $j$ is contained in at most $b$ sets, so these paths have contribution at most $b\alpha^{2l}$. Similarly, there are at most $bd$ paths of the latter variety, and each has contribution $\alpha^{2l}$. Hence,
$\Upsilon^{*}\leq\Upsilon_{G+S}({v_{s}})(2k+2)\leq\mathcal{P}_{G+S}({v_{j}},{v_{s}})(2k+2)\leq\alpha^{2l}(b+bd)(2ck+2)$
We now claim that to achieve the desired hardness bound we need only set
$l>\frac{2}{3\varepsilon}\big{[}-\log_{\alpha}(b+bd)-\log_{\alpha}(2ck+2)+\frac{4}{3}-\varepsilon\big{]}$
In this case, simple manipulations reveal that
$\frac{3\varepsilon l}{2}$ $>-\log_{\alpha}(b+bd)-\log_{\alpha}(2ck+2)+\frac{4}{3}-\varepsilon$
$\alpha^{\frac{3\varepsilon l}{2}}(b+bd)(2ck+2)$ $<\alpha^{\frac{4}{3}}\alpha^{-\varepsilon}$
$\alpha^{2l}\alpha^{\frac{3\varepsilon l}{2}}(b+bd)(2ck+2)$ $<\alpha^{2l}\alpha^{\frac{4}{3}}\alpha^{-\varepsilon}$
$\alpha^{2l}(b+bd)(2ck+2)$ $<\alpha^{2l}\alpha^{\frac{-3\varepsilon l}{2}}\alpha^{\frac{4}{3}}\alpha^{-\varepsilon}$
$\Upsilon^{*}$ $<(\alpha^{\frac{3l}{2}+1})^{\frac{4}{3}-\varepsilon}$
Consequently, any algorithm which produces reach at least $(\Upsilon^{*})^{\frac{4}{3}-\varepsilon}$ using at most $ck$ edges can also distinguish between yes- and no-instances of Gap Set Cover. This completes the proof. ∎
8 Conclusion
[…]
Interesting questions remain. Regarding our constant witness technique, we ask whether a logarithmic approximation exists for the generalization of Hitting Set in which sets are defined over bounded-size combinations of $m$ elements, i.e, the sets are subsets of $\binom{m}{c}$ for some $c\geq 1$, and the task is to find a minimum size subset of $[m]$ which contains as a subset an entire combination for each set. Also, it remains open to study Broadcast Improvement and Reach Improvement under non-uniform values of the edge propagation probability $\alpha$, or in the case of restrictions on the structure of the input graph. Lastly and most broadly, we suggest that studying algorithmic approaches to network design under stochastic models of information flow is a rich and relatively unexplored topic. Possible directions include the optimization of other metrics, e.g., the access centrality defined by [2], and the consideration of other propagation models, e.g., linear threshold [19].
Acknowledgements
This work was supported in part by the National Science Foundation under award IIS-1956286 to Blair D. Sullivan, under Grant #2127309 to the Computing Research Association for the CIFellows 2021 Project, and awards CCF-2008688 and CCF-2047288 to Aditya Bhaskara.